### v1

# 4D Scaffold Gaussian Splatting for Efficient Dynamic Scene Reconstruction

Abstract
Existing 4D Gaussian methods for dynamic scene reconstruction offer high visual fidelity and fast rendering. However, these methods suffer from excessive memory and storage demands, which limits their practical deployment. This paper proposes a 4D anchor-based framework that retains visual quality and rendering speed of 4D Gaussians while significantly reducing storage costs. Our method extends 3D scaffolding to 4D space, and leverages sparse 4D grid-aligned anchors with compressed feature vectors. Each anchor models a set of neural 4D Gaussians, each of which represent a local spatiotemporal region. In addition, we introduce a temporal coverage-aware anchor growing strategy to effectively assign additional anchors to under-reconstructed dynamic regions. Our method adjusts the accumulated gradients based on Gaussians’ temporal coverage, improving reconstruction quality in dynamic regions. To reduce the number of anchors, we further present enhanced formulations of neural 4D Gaussians. These include the neural velocity, and the temporal opacity derived from a generalized Gaussian distribution. Experimental results demonstrate that our method achieves state-of-the-art visual quality and $97.8\%$ storage reduction over 4DGS.
1 Introduction
Reconstructing dynamic scenes from multi-view videos has received significant attention due to its wide applications in augmented reality and virtual reality. Beyond methods based on neural radiance fields (NeRFs) [13, 12, 21, 28, 1, 24], 3D Gaussian Splatting (3DGS) [8] has become a major approach for dynamic scene reconstruction, due to its ability to render high-quality novel views in real-time. The two main categories of this approach are 1) modeling temporal changes as deformations of 3D Gaussians [34, 31, 2] and 2) employing 4D Gaussians to approximate a scene’s 4D volumes [35, 14].
Figure 1: Teaser. Our method achieves superior image quality, high rendering speed, and low storage with single-frame COLMAP [25] points, surpassing other 4D Gaussian methods.
Directly optimizing 4D Gaussians offers faster rendering and higher visual quality than deformation-based methods. Gaussians defined in 4D space can represent sudden changes in the scene, without computing numerous deformations of Gaussians. However, these advantages come at a storage cost: 4D Gaussians demand substantial storage, often exceeding 6GB for 10-second videos. The large number of 4D Gaussians, coupled with extensive parameter counts for 4D properties, hinders scaling this approach for real-world applications. We aim to reduce the high storage demands of 4D Gaussians, by addressing excessively repetitive patterns in dynamic scenes across space and time.
In this paper, we introduce a 4D anchor-based framework for dynamic scene reconstruction, a method that retains the visual quality and the fast rendering speed of 4D Gaussians with significantly reduced storage. In our method, dynamic scenes are represented through structured anchors aligned with a sparse 4D grid. Each anchor holds a compressed feature vector that models a set of neural 4D Gaussians, each representing a local spatiotemporal region. The properties of neural Gaussians are dynamically generated from the anchor feature vector via shared MLPs.
Our method is a non-trivial extension of 3D scaffolding [17] to 4D space, as it should carefully consider how to effectively capture dynamic regions with fewer anchors. Previous 3D scaffolding accumulates gradients of Gaussians to create new anchors, without accounting for temporal coverage. This penalizes dynamic regions that appear shortly, in contrast to background regions that are present in most frames and thus accumulate larger gradients.
We present a temporal coverage-aware anchor growing strategy to allocate new anchors to under-reconstructed dynamic regions. Our approach adjusts the accumulated gradients based on each Gaussian’s temporal coverage: Gaussians with shorter coverage are adjusted to have larger gradients. This modification prevents under-reconstructed dynamic regions from being penalized for their short appearance periods. While the proposed anchor growing operation improves the quality of reconstructed dynamic regions, it also leads to a substantial increase in the number of anchors, and associated storage costs. Such increase originates from the limited capability of generated 4D Gaussians to sufficiently capture real-world dynamics.
To this end, we propose two formulations that enhance our neural Gaussians to better represent actual scene elements. First, we model the time-varying positions of Gaussians through our neural velocity, allowing them to efficiently represent dynamics as a series of piecewise linear movements. Second, we modify our temporal opacity based on a generalized Gaussian function. This modification captures sudden changes in dynamic scenes more effectively, whereas the temporal opacity of previous methods [10, 14] lack this capability. These formulations enhance the capability of Gaussians to capture dynamics in local regions, reducing the number of anchors required for high quality.
Our 4D anchor-based reconstruction pipeline effectively addresses redundancy present in dynamic scenes through compression. As shown in Figure 1, our method achieves state-of-the-art visual quality and rendering speed, while reducing storage demands by $93.8\%$ compared to 4DGS [36].
2 Related work
This section presents a comprehensive review of scene representations, focusing on three aspects. First, we examine approaches for efficient 3D Gaussians (Sec. 2.1). Then, we investigate methods for both deforming 3D Gaussians and utilizing 4D Gaussians in dynamic scene reconstruction (Sec. 2.2). Lastly, we review various approaches employing neural features for scene reconstruction (Sec. 2.3).
Figure 2: Overview of 4D anchor-based method. We begin with a sparse 4D grid of anchors, each defined by a unique 4D spatiotemporal position $p$ and learnable offsets. Shared MLPs utilize these anchors to generate neural 4D Gaussians, capturing dynamic appearance changes with view direction $\mathbf{d}$. These 4D Gaussians are then projected to 3D Gaussians at specific timesteps for rendering, using a 3D Gaussian splatting pipeline to produce the final rendered image. We omit the $z$-axis for simplicity.
2.1 Efficient 3D Gaussians
In recent years, 3D Gaussian Splatting (3DGS) has attracted significant attention for achieving real-time rendering by representing scenes with 3D Gaussian primitives and introducing a tile-based rasterizer. For photorealistic rendering quality, 3DGS requires a significant number of 3D Gaussians, which increases storage costs. To address this problem, [6, 7, 29] remove unnecessary Gaussians that do not harm rendering quality, [16, 19, 20] replace Gaussians or their parameters with efficient representations, and [19, 4, 33, 18] compress Gaussian parameters using existing compression techniques such as entropy coding and image compression. In this paper, we address eliminating excess Gaussians and improving efficiency in dynamic scenes with 4D Gaussians, which have been relatively unexplored compared to 3D Gaussians.
2.2 Dynamic 3D Gaussians
Two main approaches are proposed to extend 3DGS [8] into dynamic scene reconstruction. The first approach involves deforming 3D Gaussians along with temporal changes [34, 31, 2, 26, 15]. These deformable 3D Gaussians offer the advantage of compact storage requirements but exhibit relatively slow rendering speeds and low visual quality. In contrast, the other approach directly employs 4D Gaussians in the spatio-temporal domain [35, 14]. These approaches demonstrate superior visual quality and faster rendering speeds, but suffer from higher storage requirements. To address this, our method efficiently models temporal opacity and predicts Gaussian properties through neural features.
2.3 Feature-based neural rendering
Recently, a growing trend in scene reconstruction has been to integrate neural features as additional inputs to enhance model performance. For instance, there have been attempts to extract features from source views and utilize them for novel view synthesis, enabling few-view reconstruction [37, 3] or view interpolation through transformers [30, 22, 27]. In 3DGS studies, [5] uses a hash grid instead of per-Gaussian SH coefficients, [14] renders features and decodes them into RGB colors via shallow MLP. [32] generates Gaussian attributes from a multi-level tri-plane. [16] predicts the attributes of local 3D Gaussians from anchor features. Our work focuses on efficient dynamic representations, which are relatively underexplored. We introduce a 4D anchor-based framework that includes dynamic neural velocity and temporal opacity derived from a generalized Gaussian distribution, considering real-world dynamics.
3 Method
We reconstruct dynamic 3D scenes with 4D Gaussians. Especially, our method builds 4D anchors which specify nearby 4D Gaussians (Sec. 3.1). We encourage anchors to be closely placed at under-reconstructed dynamic regions (Sec. 3.2). Furthermore, we design each 4D Gaussian to move along a line segment, and define the temporal opacity function which fits piecewise persistent period (Sec. 3.3).
3.1 Overview
The overview of our method is illustrated in Figure 2. Our method begins with a set of sparse 4D anchor points, each of which has a unique, grid-aligned spatiotemporal 4D position $\mathbf{p}\in\mathbb{R}^{4}$ with a feature vector $\mathbf{f}\in\mathbb{R^{C}}$. We leverage shared MLPs to produce $K$ neural 4D Gaussians from these anchor features. Corresponding 3D Gaussians are computed from these 4D Gaussians to render a frame at timestep $t$. We describe details of our method in the following paragraphs.
Initializing 4D anchors.   Similar with 3D Scaffold-GS [17], we utilize static point clouds to initialize 4D anchor points. We first obtain the static point cloud from multi-view frames at a certain timestep $t$ using Structure-from-Motion (SfM) [25]. The 4D positions of anchors are initialized from a set of voxels $V\subset\mathbb{R}^{N\times 3}$ obtained from this point cloud:
$p_{v}=(x_{v},y_{v},z_{v},t),\ \ \forall v\in V,$ (1)
where ($x_{v}$, $y_{v}$, $z_{v}$) is the center coordinates of the voxel $v$. Each anchor point also accompanies two learnable parameters: a feature vector $\mathbf{f}$, and 4D offsets $\mathbf{o}\in R^{K\times 3}$ for determining positions of $K$ neural 4D Gaussians.
Beginning with 4D anchors initialized with the static point cloud, our method gradually increases anchors to dynamic regions. This is accomplished by our anchor growing strategy, which will be described in Sec. 3.2.
Neural 4D Gaussians.   We leverage shared MLPs and learnable parameters of the 4D anchors to generate $K$ neural 4D Gaussians from each anchor. Specifically, shared MLPs take the anchor feature $\mathbf{f}$ and yield properties of $K$ neural Gaussians, which include time-invariant opacity $\alpha\in\mathbb{R}$, quaternion $\mathbf{q}\in\mathbb{R}^{4}$ and scaling $\mathbf{s}\in\mathbb{R}^{3}$ for the covariance matrix, view-dependent color $\mathbf{c}\in\mathbb{R}^{3}$. Our shared MLPs also produce temporal scale $\sigma\in\mathbb{R}$ to compute our temporal opacity, and the neural velocity $\mathbf{v}\in\mathbb{R}^{3}$. Note that the color MLP additionally takes the view direction $\mathbf{d}\in\mathbb{R}^{3}$ as inputs for view-dependent modeling.
Rendering neural Gaussians.   To render the neural 4D Gaussians, we compute 3D Gaussian parameters at time $t$ from our time-varying properties. For the $k$-th Gaussian of the anchor $\mathbf{p}$, the opacity $\alpha_{k}(t)$ and the center $\mu_{k}(t)$ of the corresponding 3D Gaussian at time $t$ are derived as
$\mu_{k}(t)=\mu_{k,1:3}+h(t,\mu_{k,4},\mathbf{v}),$ (2)
$\alpha_{k}(t)=\alpha_{k}\cdot g(t,\mu_{k,4},\sigma_{k}),$ (3)
where $\mu_{k}=\mathbf{p}+\mathbf{o}_{k}$ is the coordinates of the 4D Gaussian. $h(\cdot)$, $g(\cdot)$ model time-varying values of the positions and the opacity with the neural velocity and temporal opacity function, which will be further described in Sec. 3.3.
After deriving 3D Gaussians at time $t$, we utilize the existing efficient 3D Gaussian splatting pipeline [8] to render the frames. Same as 3D Scaffold-GS [17], Gaussians within the view frustum and having opacity higher than the threshold are passed to the renderer.
Figure 3: Time-variant Gaussian position calculation. The 3D Gaussian position at rendering time is computed as the sum of the anchor position, the 4D Gaussian offset, and the temporal linear movement. The temporal linear movement is the product of the neural velocity and the time difference, allowing the Gaussian to traverse 3D space over time.
3.2 Temporal coverage-aware anchor growing
Growing new anchors to under-reconstructed dynamic regions is a critical factor for high-quality results. The direct application of previous anchor growing to the dynamic reconstruction fails to identify such regions, as it neglects the temporal coverage of the Gaussians.
To this end, we propose a temporal coverage-aware anchor growing operation. The core of our anchor growing is the computation of the accumulated gradients of each 4D Gaussian $\nabla_{g}$. We formulate $\nabla_{g}$ as a weighted sum of the gradient over $N$ iterations:
$\nabla_{g}=\frac{\sum^{N}w(t,\sigma)\|\nabla_{\text{2D}}\|}{\sum^{N}w(t,\sigma)},$ (4)
where $\nabla_{\text{2D}}$ is the 2D position gradient of the Gaussian. We define the weight term $w$ as a function of $\alpha(t)$ and $\sigma$, with $\gamma$ as a hyperparameter:
$w(t,\sigma)=\alpha(t)(1/\sigma)^{\gamma}.$ (5)
We collect the accumulated gradients of all Gaussians and then voxelize them. New anchors are placed at the centers of the voxels having gradients higher than a threshold.
In contrast, the previous anchor growing strategy [17] designed for static scenes gathers the gradients as a mean over $N$ iterations:
$\nabla_{g}=\frac{\sum^{N}\|\nabla_{\text{2D}}\|}{N}.$ (6)
With this strategy, dynamic regions appearing in a short period will have lower $\nabla_{g}$, as these regions will be penalized by the denominator $N$ regardless of their actual errors.
The first difference of our approach compared to the previous anchor growing strategy [17] is that the gradients are accumulated only when the Gaussian is activated. Our approach accurately gathers the gradients of 4D Gaussians placed at the regions where dynamic scene elements only appear in a short period. The second difference is that the temporal coverage $\sigma$ directly influences to the anchor growing operation. This makes regions having a short temporal coverage easier to grow. By modifying the accumulated gradients as a weighted sum, our approach effectively places new anchors to under-reconstructed dynamic regions.
3.3 Modeling 4D Gaussians for compact anchors
To represent dynamic scenes with fewer anchors, each neural Gaussian should be able to model the scene element in a sufficient time range. This necessitates neural 4D Gaussians to accurately model the real-world dynamics.
We introduce two key formulations of 4D Gaussians that improve the capability of Gaussians of modeling real-world dynamics. We model local movements through the neural velocity, and we modify the temporal opacity based on a generalized Gaussian distribution.
Neural velocity.   In computer graphics, one common way to model the motion of an object is through a set of piecewise linear segments. Based on this concept, we propose the neural velocity to model the motions of scene elements with a set of 4D Gaussians having linear movements (see Figure 3). As described in Equation (2), the time-varying 3D position of the 3D Gaussian $\mu_{k}(t)\in\mathbb{R}^{3}$ is determined by the time-varying function $h(\cdot)$. We formulate the $h(\cdot)$ as the linear movement based on the neural velocity $\mathbf{v}$:
$h(t,\mu_{k,4},\mathbf{v})=(t-\mu_{k,4})\mathbf{v}.$ (7)
Note that our neural velocity simplifies the temporal slicing of 4D positions in 4DGS [10], without requiring excessive parameters for 4D covariances.
Despite its simplicity, our neural velocity enables each 4D Gaussian to cover dynamic movements of the elements in a sufficient time range. Such capability reduces the number of neural 4D Gaussians and anchors as a result.
Figure 4: Illustration of the modified temporal opacity. To reconstruct the sudden and continuous appearance of an object, the previous univariate Gaussian model requires the sum of multiple Gaussians, while the generalized Gaussian achieves the same expressiveness with only a single Gaussian.
Modified temporal opacity.   Previous 4D Gaussian methods [10, 14] adopt the univariate Gaussian function as temporal opacity to make Gaussians appear within a specific time range. However, this univariate Gaussian lacks the ability to model sudden changes in real-world scene elements, resulting in excessive Gaussians to model a single scene element. This issue is also discussed in the concurrent work [11], which proposes a simplified Gaussian mixture model as a solution.
Therefore, we propose to modify the temporal opacity function to remove such redundant Gaussians (see Figure 4). We reformulate the temporal opacity function based on a generalized Gaussian distribution. This offers steeper derivatives at the beginning and the end, making it better suited to fit piecewise persistent periods. Our time-varying opacity $g(\cdot)$ in Equation (3) is formulated as
$g(t,\mu_{k,4},\sigma_{k})=e^{-(|t-\mu_{k,4}|/\sigma_{k})^{\beta}}.$ (8)
This formulation greatly reduces the number of 4D Gaussians and anchors to represent dynamic scenes while preserving the visual quality. In practice, we set the inverse sigma as a learnable parameter, as it leads to more stable training. We also model $\beta=2\beta^{\prime}$ and $\beta^{\prime}$ as a hyperparameter, which removes the ||$\cdot$|| ⋅ | in the above equation.
3.4 Training and implementation details
Table 1: Quantitative results on N3DV dataset. 1 only includes the flame_salmon for storage and FPS [2]. 2 divides the scene into 50 frames and trains each using a separate model. 3 uses COLMAP [25] for all 300 frames.
[…]
K-Planes [24] 1 31.22 0.947 0.090 309 MB 0.13
MixVoxels-L [28] 1 30.81 0.933 0.095 512 MB 0.93
HyperReel [1] 1 2 31.10 0.927 0.099 1362 MB 1.04
4DGaussian [31] 30.71 0.935 0.056 59 MB 51.9
E-D3DGS [2] 31.42 0.945 0.037 137 MB 19.9
STG [14] 2 3 31.96 0.948 0.046 173 MB 125.9
4DGS [35] 32.14 0.947 0.047 6194 MB 61.4
Ours 32.03 0.947 0.041 134 MB 129.9
Training objective.   Both learnable anchor parameters and MLP weights are jointly optimized through the rendering loss. We employ $\mathcal{L}_{1}$ with SSIM loss $\mathcal{L}_{\text{SSIM}}$ and volume regularization $\mathcal{L}_{\text{vol}}$ following the 3D Scaffold-GS [16]. The full training objective with weighting coefficients $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{vol}}=0.01$ is
$\mathcal{L}=(1-\lambda_{\text{SSIM}})\mathcal{L}_{1}+\lambda_{\text{SSIM}}\mathcal{L}_{\text{SSIM}}+\lambda_{\text{vol}}\mathcal{L}_{\text{vol}}.$ (9)
Implementation details.   Similar to 4DGS, our model is trained for 120K iterations with a single batch, which takes approximately 4 hours on a single NVIDIA A6000 GPU. We set $\beta=2$ and $\gamma=1$ for our modified temporal opacity model and temporal coverage-aware densification. For 4D voxel grid size, We use 0.001 as spatial grid size and use frame interval time as temporal grid size for all scenes. After training, we prune the unnecessary anchor points of which all Gaussians have opacity near zero. To improve rendering speed, we cache the time-invariant and view-invariant outputs of MLPs. We provide more implementation details in the supplements.
4 Experiments
In this section, we first evaluate our method through the comparisons with several state-of-the-art baselines for dynamic scene reconstruction. Then we conduct ablations and analysis to explore the effectiveness of our main features. Our code will be made publicly available.
Datasets.   We evaluate our method on two representative real-world datasets: Neural 3D Video dataset (N3DV) [12] and Technicolor dataset [23].
  • Neural 3D Video dataset (N3DV) comprises 17 to 21 synchronized videos of six scenes, with each video containing 300 frames. Following previous works, we downsample the resolution to $1352\times 1014$ and use the first camera as the test view. We exclude cam13 for the coffee_martini scene due to synchronization issues.
  • Technicolor Dataset contains 16 synchronized videos across five scenes, with each video comprising 50 frames. We retain the original resolution of $2048\times 1088$ and use cam10 as the test view.
Baselines.   We choose the following state-of-the-art competitors: 4DGaussian [31] and E-D3DGS [2] representing deformation-based methods, and 4DGS [35] and STG [14] representing 4D Gaussians. We reproduce them using their official code to report their performance. We also include representatives of NeRF-based methods: K-Planes [24], MixVoxels [28], and HyperReel [1]. Results of these methods are brought from [2].
Metrics.   To assess the reconstruction quality, we measure peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and LPIPS [38] of the rendered images. We employ AlexNet [9] to compute LPIPS following the previous papers. To evaluate storage efficiency, we calculate the total size of output files, including MLP weights and parameters of Gaussians or anchors. We measure frames-per-second (FPS) by recording the wall-clock rendering time on the same machine, with a single NVIDIA A6000 GPU.
Table 2: Quantitative results on Technicolor dataset. 1 refers to the values reported in [14]. 2 only includes the Painter for storage and FPS [2]. 3 uses COLMAP [25] for all 50 frames.
[…]
DyNeRF [12] 1 31.80 - 0.140 30 MB 0.02
HyperReel [1] 2 32.32 0.899 0.118 289 MB 0.91
4DGaussian [31] 29.62 0.844 0.176 51 MB 34.8
E-D3DGS [2] 33.24 0.907 0.100 77 MB 36.5
STG [14] 2 3 33.40 0.918 0.082 55 MB 112.7
4DGS [35] 33.35 0.910 0.095 9106 MB 43.3
Ours 33.97 0.923 0.074 98 MB 124.5
Figure 5: Qualitative comparisons on the Neural 3D Video dataset. This figure provides qualitative comparisons on the Neural 3D Video dataset. White boxes indicate under-reconstructed areas. Our method achieves high fidelity in both dynamic and static regions.
4.1 Reconstruction quality
N3DV.   Table 1 presents the quantitative results on the N3DV dataset. Our method outperforms all NeRF baselines in both visual quality and rendering speed. It also achieves competitive results in visual quality compared to state-of-the-art Gaussian Splatting-based methods, and offers the fastest rendering speed across all methods. Remarkably, our method significantly reduces storage costs by 97.8×97.8$\times$97.8 × compared to 4DGS [35], while achieving comparable or better performance in visual quality and rendering speed. It is worth noting that STG requires COLMAP for all frames, and train 6 models to obtain its results. We provide further comparisons with STG in the supplements.
To further assess the performance, we deliver the qualitative comparisons with Gaussian Splatting baselines in Figure 5. Our method successfully reconstructs both static and dynamic regions. In contrast, other baselines often struggle to reconstruct dynamic regions (see the boxed regions). Deformation-based methods (4DGaussians, E-D3DGS) tend to produce blurry results, and the results of 4D Gaussian-based methods (4DGS, STG) often contain artifacts in dynamic regions. By effectively modeling dynamic regions with fewer anchors, our method delivers high-quality results in an efficient manner. Please refer to the supplements for more visual comparisons, including uncropped results and videos.
Figure 6: Qualitative comparisons on the Technicolor dataset. This figure provides qualitative comparisons on the Technicolor dataset. White boxes indicate under-reconstructed areas. Our method achieves high fidelity in both dynamic and static regions.
Technicolor.   We deliver the quantitative comparisons on the Technicolor dataset in Table 2. Our method surpasses all baselines in visual quality and rendering speed. We also achieve 85.1×85.1$\times$85.1 × storage reduction and $\times2$ improvement in FPS, compared to 4DGS. Although deformation-based methods offer compact storage usage, these methods require expensive computations for Gaussian deformations. This leads to the low FPS of these methods.
Figure 6 presents the qualitative comparisons with Gaussian-based methods on the Technicolor dataset. Our method produces sharp and high-fidelity rendering results in both static and dynamic regions. On the other hand, 4DGaussians, E-D3DGS and STG show blurry results in dynamic regions, such as the boxed part in the Theater scene. E-D3DGS and STG often fail to reconstruct several details, e.g. the eye in the Theater scene.
4.2 Analysis and ablation study
To provide deeper insights of our main features, we conduct the ablations and analysis on the N3DV dataset. Please refer to the supplements for more detailed analysis and ablations on each component.
Effects of each feature on quality and storage.   We verify the effectiveness of our proposals through the ablations. We begin with the baseline model (Baseline) that does not employ none of our main components. Then we gradually add each component to the baseline and report its results.
As shown in Table 3, our model with all components achieves the best results in visual quality with reasonable storage. Adding the temporal coverage-aware anchor growing operation improves visual quality but increases the anchors. The neural velocity reduces the number of anchors but brings degradation in visual quality. Finally, modifying the temporal opacity achieves the highest visual quality with fewer anchors. These results support that our anchor growing strategy successfully allocates new anchors to under-reconstructed dynamic regions, improving visual quality. Appropriate modeling of neural Gaussians reduces the number of anchors without sacrificing performance.
Table 3: Ablation results on the main components. The abbreviations are as follows: “TA” refers to the temporal coverage-aware anchor growing, “NV” to the neural velocity and “MT” to the modified temporal opacity.
[…]
31.20 0.053 207K 83 MB
31.74 0.043  [^8] K 301 MB
31.11 0.042 576K 195 MB
31.82 0.043 793k 268 MB
32.03 0.041 440K 149 MB
Figure 7: Anchor growing analysis. Top-to-bottom: accumulated gradients at 5000th training iteration, final 4D anchors at 106 frame, and the rendering results. Our anchor growing effectively accumulates gradients in under-reconstructed dynamic regions.
Understanding anchor growing.   To provide an in-depth understanding of the proposed anchor growing strategy, we present visual comparisons with previous anchor growing operations. We visualize the accumulated gradients $\nabla_{g}$ of both methods during training. We also compare resulting anchors and rendered images after training.
Figure 7 presents visual comparisons with the previous anchor growing operation. The previous method fails to accurately collect gradients in dynamic regions due to their short appearing periods. This leads to a lack of anchors in under-reconstructed regions, resulting in degraded visual quality. In contrast, our method successfully accumulates higher gradients in these dynamic regions (see the red-boxed parts in Figure 7). Consequently, the anchors generated by our method more accurately represent dynamic regions at each timestep. These results demonstrate that our anchor-growing strategy effectively captures under-reconstructed dynamic regions, playing a critical role in achieving high-quality reconstruction.
Figure 8: Temporal opacity analysis. We visualize Gaussians that contributing to the rendering of a sample pixel in the sear_steak scene from the N3DV dataset. For clarity, we filter out Gaussians distant from the frying pan based on depth and image-plane probability. Our modified temporal opacity better aligns with real-world dynamics, allowing a single Gaussian to represent the element.
Understanding temporal opacity.   We explore the effects of the temporal opacity by observing actual distributions. Specifically, we choose an example pixel in sear_steak, which contains an object that appears in a certain period (the frying pan). We visualize Gaussians that primarily contribute to render this pixel. We compare our temporal opacity with the previous one based on a univariate Gaussian function.
As shown in Figure 8, the resulting Gaussian with our modified temporal opacity better fits the actual distributions of the object. This enables our method to represent a scene element with a single Gaussian. On the other hand, previous univariate Gaussian is not well-aligned to real-world dynamics, resulting in multiple Gaussians representing the same element. As a result, our temporal opacity reduces the number of anchors while retaining visual quality, which leads to compact storage usage.
5 Conclusion
We proposed the 4D anchor-based method for dynamic scene reconstruction. Our anchor growing and neural 4D Gaussians modeling capture dynamic regions with an efficient number of anchors. Experimental results support the validity of our method, achieving compelling visual quality and FPS while significantly reducing storage consumption.
Limitations and future work.   Since our method is currently designed for multi-view video datasets, applying it on monocular videos can introduce additional challenges. Same as other methods, our method still suffer from reconstructing elements that appear very shortly (1 or 2 frames). Resolving this challenge can be an interesting future work.
Supplementary Material
Figure S1: Comparison between 300-frame STG and Ours. The single STG model trained on 300 frames of the N3DV dataset shows lower quality in the dynamic regions compared to Ours. The white boxes highlight under-reconstruction areas.
Table S1: Comparison with STG on the N3DV dataset. Our model outperforms the average of six STGs trained on 50 frames and a single STG trained on 300 frames. $\ast$ Note that STG uses a batch size of 2 while our model uses a batch size of 1.
[…]
STG 6 (50 frames) 6$\times$30K 31.96 0.046
1 (300 frames) 60K 31.25 0.053
Ours 1 (300 frames) 120K 32.03 0.041
Appendix A Additional comparison with STG
As mentioned in the main paper, it is difficult to fairly compare with Spacetime Gaussians (STG) [14] due to the different experimental setups. Unlike our method, STG utilizes per-frame COLMAP to obtain the initial points and divide long video sequences into 50-frame chunks, training separate models for each chunk. This results in six separate models to report its results on N3DV.
We additionally deliver the results of STG on N3DV, under the same experimental setup in Figure S1 and Table S1. While STG is effective on shorter video sequences, it often fails to accurately reconstruct dynamic regions on longer sequences. On the other hand, our method effectively represents dynamic regions on longer sequences. As a result, our method achieves better visual quality than STG under the same experimental setups.
Appendix B Modified temporal opacity with 4DGS.
We further validate the effectiveness of our modified temporal opacity by applying it to 4DGS [35]. Consistent with the main paper, we visualize temporal opacity values at the same points within sear_steak scene of the N3DV dataset. $\beta=8$ is used for the generalized Gaussian equation.
As shown in Figure S2, the previous univariate Gaussian function struggles to represent elements that suddenly appear and disappear. This limitation results in five Gaussians to model the same point. On the other hand, our modified temporal opacity effectively represents these dynamic changes, enabling the representation with only one Gaussian. Consequently, 4DGS with the modified temporal opacity achieves a storage reduction from 4.74 GB to 3.82 GB.
Figure S2: Impact of modified temporal opacity with 4DGS. We apply our modified temporal opacity to 4DGS and analyze its behavior on the sear_steak scene from N3DV. The total temporal opacity distribution for the frying pan (blue point) is low at $t=20$ and $t=90$ while peaking at $t=60$, with rapid transitions in between. Each plot represents the temporal opacity of an individual Gaussian. Unlike conventional methods relying on Gaussian mixtures, our generalized Gaussian effectively captures this behavior using a single Gaussian component.
Appendix C Additional ablation study
Modeling 4D Gaussians without anchors.   To analyze the effectiveness of our anchor-based reconstruction pipeline, we evaluate the performance of our model without the scaffold structure. Specifically, we replace all MLP outputs with explicit learnable Gaussian parameters. As shown in Table S2, removing the scaffold structure leads to a degradation in visual quality. The structured anchor representations offer more stable training and anchor growing, which significantly improves the performance of our method.
Polynomial Gaussian trajectory.   Our neural velocity represents the local dynamics of Gaussians as linear movements. Despite its simplicity, these linear movements sufficiently represent dynamics as 3D scene flows. We explore the effectiveness of the neural velocity by replacing it with a polynomial trajectory proposed in the previous method [14]. We use $n_{p}=3,n_{q}=1$ as they use for their full model. The velocity MLP outputs coefficients of the polynomial equation instead of the neural 3D velocity. As shown in Table S2, this approach results in decreased performance and increased storage. Since the polynomial trajectory requires 12 parameters per Gaussian, a total of 120 parameters ($K=10$) as MLP outputs, it often complicates the training of our lightweight MLP and leads to degradation.
Table S2: Additional model ablations. Quantitative results on different model designs on the N3DV dataset. Our full method utilizes scaffold structure and neural velocity.
[…]
velocity 30.11 0.060 393 MB
polynomial 31.68 0.046 323 MB
velocity 32.03 0.041 134 MB
Appendix D Effects of hyperparameters
Figure S3: Effect of $\gamma$. (upper) As the value of $\gamma$ increases, the quality of dynamic regions improves, whereas the quality of static regions declines. (lower) The number of anchor points changes with $\gamma$.
Figure S4: Effect of $\beta$. The graph shows how $\beta$ affects the number of anchor points across two scenes, with an increase in $\beta$ leading to a reduction in the number of anchor points.
Figure S5: Visualization of Anchor Growing. The blue dots on the left represent the initial anchor points, which are obtained exclusively at $t=0$. For a specific frame at $t=194$, new anchor points grow as iterations progress, shown as green dots on the right. These new anchor points are primarily generated in under-reconstructed dynamic regions.
We investigate the impact of hyperparameters on the performance and their controllability. cook_spinach and flame_salmon scenes from N3DV are selected for analysis.
Temporal coverage-aware anchor growing.   The parameter $\gamma$ in Equation (5) controls the influence of temporal scale on densification. As $\gamma$ increases, densification becomes more concentrated in dynamic regions, improving reconstruction quality in these areas while reducing quality in static regions, as demonstrated in Figure S3. In scenes with distant background regions, such as flame_salmon, larger $\gamma$ values result in less densification in the static background, leading to a decline in reconstruction quality. Additionally, excessively high $\gamma$ values result in over-densification in dynamic regions, increasing the number of anchors. We empirically observe that $\gamma=1$ provides a balanced reconstruction between static and dynamic regions.
Modified temporal opacity.   The parameter $\beta$ determines the steepness of our generalized Gaussian curve. Higher values of $\beta$ result in steeper slopes, enabling shapes to be represented with fewer Gaussian components compared to a typical mixture of univariate Gaussians. As shown in Figure S4, increasing $\beta$ reduces the number of required anchor points while maintaining image quality. However, excessively high $\beta$ values (e.g., above 8) cause instability during training.
Appendix E Visualization of anchor growing
We visualize the anchor growing process in Figure S5. During the early stages of training, the temporal scale of 4D Gaussians generated from the initial anchors increases, enabling rapid reconstruction of static regions. As training progresses, anchors are created in the dynamic regions at each time step $t$ in under-reconstructed regions, improving image quality in these areas.
Appendix F Additional method detail
Network architecture.   In this section, we describe the detailed architecture of the MLPs utilized in our model. Each anchor feature is represented as a 32-dimensional vector, which is passed through the shared MLPs to compute the properties of $K$ neural 4D Gaussians. Each MLP consists of a 2-layer architecture with a hidden layer of width 32, matching the size of the feature dimension. The hidden layer uses a ReLU activation function. We employ a total of four MLPs, referred to as the Opacity MLP, Shape MLP, Color MLP, and Velocity MLP.
The Opacity MLP outputs the time-invariant opacity $\alpha$ of the Gaussians. The Shape MLP produces the quaternion q and scaling s for covariance calculation, along with the temporal scale $\sigma$. The Color MLP takes the direction d concatenated with the anchor feature as input and outputs the view-dependent color c. The Velocity MLP computes the 3D velocity of the Gaussians. $\tanh$ for opacity $\alpha$, $\exp$ for scaling s and temporal scale $\sigma$, and sigmoid for color c are applied as activation function.
Implmentation detail.   The temporal grid size is set to 0.0333 for the N3DV dataset and 0.02 for the Technicolor dataset. For initial points, we downsample it to fewer than 100,000 points. All other hyperparameters, including the learning rate and learning decay schedule, follow those of [16]. Components such as the feature bank, LOD, and appearance features in [16] are excluded.
Appendix G More results
We report the quantitative results in Table S3-S4 and qualitative results in Figure S6-S7 of each scene of the N3DV and the Technicolor datasets, respectively. We also provide video quality comparisons in HTML format. Please check the index.html file.
Table S3: Per-scene quantitative results on the N3DV dataset.
Model Metric coffee_martini cook_spinach cut_roasted_beef
[…]
4DGaussians 28.44 0.919 0.060 33.10 0.953 0.042 33.12 0.954 0.044
E-D3DGS 29.10 0.931 0.042 32.96 0.956 0.034 33.57 0.958 0.034
4DGS 28.63 0.918 0.071 33.54 0.956 0.039 34.18 0.959 0.038
STG 28.26 0.915 0.074 33.07 0.958 0.038 33.38 0.959 0.039
Ours 28.80 0.923 0.060 33.36 0.955 0.036 33.35 0.957 0.036
Model Metric flame_salmon flame_steak sear_steak
[…]
4DGaussians 28.80 0.926 0.060 33.55 0.961 0.032 34.02 0.963 0.031
E-D3DGS 29.61 0.936 0.038 33.57 0.964 0.028 33.45 0.963 0.030
4DGS 29.25 0.929 0.057 33.90 0.961 0.037 33.37 0.960 0.040
STG 29.44 0.925 0.067 33.82 0.965 0.031 33.77 0.966 0.031
Ours 29.20 0.928 0.054 33.73 0.960 0.030 33.73 0.960 0.031
Table S4: Per-scene quantitative results on the Technicolor dataset.
Model Metric Birthday Fabien Painter
[…]
4DGaussians 28.03 0.862 0.156 33.36 0.865 0.185 34.52 0.899 0.138
E-D3DGS 33.34 0.951 0.037 34.45 0.875 0.147 36.63 0.923 0.097
4DGS 31.81 0.937 0.040 35.02 0.884 0.144 35.71 0.923 0.104
STG 31.88 0.947 0.028 35.79 0.900 0.115 35.53 0.925 0.095
Ours 32.86 0.952 0.024 35.32 0.889 0.140 37.35 0.941 0.069
Model Metric Theater Train
[…]
4DGaussians 28.67 0.835 0.195 23.54 0.756 0.206
E-D3DGS 30.64 0.864 0.147 31.84 0.922 0.074
4DGS 31.94 0.875 0.143 32.25 0.932 0.046
STG 31.25 0.880 0.133 32.58 0.938 0.039
Ours 31.65 0.891 0.106 32.69 0.941 0.032
Figure S6: Qualitative comparisons on the N3DV dataset.
Figure S7: Qualitative comparisons on the Technicolor dataset.

### v2

@@
-4D Scaffold Gaussian Splatting for Efficient Dynamic Scene Reconstruction
+4D Scaffold Gaussian Splatting with Dynamic-Aware Anchor Growing for Efficient and High-Fidelity Dynamic Scene Reconstruction
 Abstract
-Existing 4D Gaussian methods for dynamic scene reconstruction offer high visual fidelity and fast rendering.
-However, these methods suffer from excessive memory and storage demands, which limits their practical deployment.
-This paper proposes a 4D anchor-based framework that retains visual quality and rendering speed of 4D Gaussians while significantly reducing storage costs.
-Our method extends 3D scaffolding to 4D space, and leverages sparse 4D grid-aligned anchors with compressed feature vectors.
-Each anchor models a set of neural 4D Gaussians, each of which represent a local spatiotemporal region.
-In addition, we introduce a temporal coverage-aware anchor growing strategy to effectively assign additional anchors to under-reconstructed dynamic regions.
-Our method adjusts the accumulated gradients based on Gaussians’ temporal coverage, improving reconstruction quality in dynamic regions.
-To reduce the number of anchors, we further present enhanced formulations of neural 4D Gaussians.
-These include the neural velocity, and the temporal opacity derived from a generalized Gaussian distribution.
-Experimental results demonstrate that our method achieves state-of-the-art visual quality and $97.8\%$ storage reduction over 4DGS.
-1 Introduction
-Reconstructing dynamic scenes from multi-view videos has received significant attention due to its wide applications in augmented reality and virtual reality.
-Beyond methods based on neural radiance fields (NeRFs) [13, 12, 21, 28, 1, 24], 3D Gaussian Splatting (3DGS) [8] has become a major approach for dynamic scene reconstruction, due to its ability to render high-quality novel views in real-time.
-The two main categories of this approach are 1) modeling temporal changes as deformations of 3D Gaussians [34, 31, 2] and 2) employing 4D Gaussians to approximate a scene’s 4D volumes [35, 14].
-Figure 1: Teaser. Our method achieves superior image quality, high rendering speed, and low storage with single-frame COLMAP [25] points, surpassing other 4D Gaussian methods.
-Directly optimizing 4D Gaussians offers faster rendering and higher visual quality than deformation-based methods.
-Gaussians defined in 4D space can represent sudden changes in the scene, without computing numerous deformations of Gaussians.
-However, these advantages come at a storage cost: 4D Gaussians demand substantial storage, often exceeding 6GB for 10-second videos.
-The large number of 4D Gaussians, coupled with extensive parameter counts for 4D properties, hinders scaling this approach for real-world applications.
-We aim to reduce the high storage demands of 4D Gaussians, by addressing excessively repetitive patterns in dynamic scenes across space and time.
-In this paper, we introduce a 4D anchor-based framework for dynamic scene reconstruction, a method that retains the visual quality and the fast rendering speed of 4D Gaussians with significantly reduced storage.
-In our method, dynamic scenes are represented through structured anchors aligned with a sparse 4D grid.
-Each anchor holds a compressed feature vector that models a set of neural 4D Gaussians, each representing a local spatiotemporal region.
-The properties of neural Gaussians are dynamically generated from the anchor feature vector via shared MLPs.
-Our method is a non-trivial extension of 3D scaffolding [17] to 4D space, as it should carefully consider how to effectively capture dynamic regions with fewer anchors.
-Previous 3D scaffolding accumulates gradients of Gaussians to create new anchors, without accounting for temporal coverage.
-This penalizes dynamic regions that appear shortly, in contrast to background regions that are present in most frames and thus accumulate larger gradients.
-We present a temporal coverage-aware anchor growing strategy to allocate new anchors to under-reconstructed dynamic regions.
-Our approach adjusts the accumulated gradients based on each Gaussian’s temporal coverage: Gaussians with shorter coverage are adjusted to have larger gradients.
-This modification prevents under-reconstructed dynamic regions from being penalized for their short appearance periods.
-While the proposed anchor growing operation improves the quality of reconstructed dynamic regions, it also leads to a substantial increase in the number of anchors, and associated storage costs.
-Such increase originates from the limited capability of generated 4D Gaussians to sufficiently capture real-world dynamics.
-To this end, we propose two formulations that enhance our neural Gaussians to better represent actual scene elements.
-First, we model the time-varying positions of Gaussians through our neural velocity, allowing them to efficiently represent dynamics as a series of piecewise linear movements.
-Second, we modify our temporal opacity based on a generalized Gaussian function.
-This modification captures sudden changes in dynamic scenes more effectively, whereas the temporal opacity of previous methods [10, 14] lack this capability.
-These formulations enhance the capability of Gaussians to capture dynamics in local regions, reducing the number of anchors required for high quality.
-Our 4D anchor-based reconstruction pipeline effectively addresses redundancy present in dynamic scenes through compression.
-As shown in Figure 1, our method achieves state-of-the-art visual quality and rendering speed, while reducing storage demands by $93.8\%$ compared to 4DGS [36].
-2 Related work
-This section presents a comprehensive review of scene representations, focusing on three aspects.
-First, we examine approaches for efficient 3D Gaussians (Sec.
-2.1).
-Then, we investigate methods for both deforming 3D Gaussians and utilizing 4D Gaussians in dynamic scene reconstruction (Sec.
-2.2).
-Lastly, we review various approaches employing neural features for scene reconstruction (Sec.
-2.3).
-Figure 2: Overview of 4D anchor-based method.
-We begin with a sparse 4D grid of anchors, each defined by a unique 4D spatiotemporal position $p$ and learnable offsets.
+Modeling dynamic scenes through 4D Gaussians offers high visual fidelity and fast rendering speeds, but comes with significant storage overhead.
+Recent approaches mitigate this cost by aggressively reducing the number of Gaussians.
+However, this inevitably removes Gaussians essential for high-quality rendering, leading to severe degradation in dynamic regions.
+In this paper, we introduce a novel 4D anchor-based framework that tackles the storage cost in different perspective.
+Rather than reducing the number of Gaussians, our method retains a sufficient quantity to accurately model dynamic contents, while compressing them into compact, grid-aligned 4D anchor features.
+Each anchor is processed by an MLP to spawn a set of neural 4D Gaussians, which represent a local spatiotemporal region.
+We design these neural 4D Gaussians to capture temporal changes with minimal parameters, making them well-suited for the MLP-based spawning.
+Moreover, we introduce a dynamic-aware anchor growing strategy to effectively assign additional anchors to under-reconstructed dynamic regions.
+Our method adjusts the accumulated gradients with Gaussians’ temporal coverage, significantly improving reconstruction quality in dynamic regions.
+Experimental results highlight that our method achieves state-of-the-art visual quality in dynamic regions, outperforming all baselines by a large margin with practical storage costs.
+Figure 1: Dynamic region quality comparison.
+(left) Comparison of dynamic region quality ($y$-axis) versus storage cost ($x$-axis) on N3DV.
+(right) Results on the flame_salmon scene from N3DV.
+Scaff-naive refers to the naive anchor-based model without our neural Gaussian design and dynamic-aware anchor growing.
+Our method significantly outperforms all baselines with efficient storage usage, while other methods either exhibit degraded quality in dynamic regions or excessive storage costs.
+Introduction
+Reconstructing dynamic scenes from multi-view videos has received significant attention due to its wide applications.
+Beyond methods based on neural radiance fields (NeRFs) (Li et al.
+2022; Wang et al.
+2022; Attal et al.
+2023), 3D Gaussian Splatting (3DGS) (Kerbl et al.
+2023) has become a major approach for dynamic scene reconstruction, due to its ability to render high-quality novel views in real-time.
+The two main categories of this approach are 1) modeling temporal changes as deformations of canonical 3D Gaussians (Yang et al.
+2023; Bae et al.
+2024; Jiawei et al.
+2024) and 2) employing 4D Gaussians to approximate a scene’s 4D volumes (Yang et al.
+2024; Li et al.
+2023; Lee et al.
+2024c, a).
+Directly optimizing 4D Gaussians offers higher visual quality and faster rendering than deformation-based methods.
+By modeling temporal changes through multiple 4D Gaussians, each covering a certain time range, they effectively capture complex dynamic regions without the heavy computation cost of deformation fields.
+However, this approach accompanies a large number of 4D Gaussians, which subsequently leads to substantial storage overhead–often exceeding 6GB for a 10-second video (see Figure 1, left).
+Several approaches have attempted to address the storage cost by reducing the number of Gaussians.
+They achieve this by more complicated motion modeling (Lee et al.
+2024b), aggressive pruning (Li et al.
+2023), or motion interpolation (Lee et al.
+2024a).
+While these methods effectively reduce storage, they also inevitably remove Gaussians in dynamic regions, sacrificing the expressiveness needed to represent complex temporal changes.
+As a result, these methods often suffer from quality degradation in dynamic components.
+In this paper, we introduce a 4D anchor-based framework that addresses the storage overhead from a different perspective.
+Instead of reducing the number of Gaussians–which is crucial for maintaining the expressiveness–our method maintains a sufficient number to render dynamic regions with high visual quality.
+This is achieved by representing dynamic scenes through structured anchor features (Lu et al.
+2024a), aligned with a sparse 4D grid.
+Each anchor holds a compact feature vector to model a local spatiotemporal region.
+This feature is processed by shared multi-layer perceptrons (MLP) to output properties of neural Gaussians.
+Considering the limited capacity of the shared shallow MLP, we propose a compact parametrization of the neural Gaussians that effectively captures temporal changes.
+This includes two key modelings: linear motion to model the time-varying 3D position, and a generalized Gaussian function to model temporal opacity, which is better suited for capturing sudden appearance changes.
+This design enables our framework to effectively represent complex dynamic trajectories through a sequence of linear segments.
+While the anchor-based scheme effectively reduces storage, a naïve extension of 3D scaffolding (Lu et al.
+2024a) fails to accurately capture dynamic regions (see Figure 1, Scaff-naive).
+The limitation arises from the anchor growing strategy used in (Lu et al.
+2024a), which is originally designed for static scenes.
+In our 4D framework, each anchor and its neural Gaussians are active only within a specific time range, while the anchor growing strategy accumulates gradients across all frames without considering this coverage.
+As a result, dynamic regions appearing in only a few frames receive less gradient signals, leading to under-reconstructed dynamic components in the final scene.
+We present a dynamic-aware anchor growing strategy to properly allocate new anchors to under-reconstructed dynamic regions.
+Our approach adjusts the accumulated gradients based on each Gaussian’s temporal coverage.
+We adjust dynamic Gaussians with shorter coverage to have larger gradient, compensating for the penalty caused by their short appearance.
+This modification encourages new anchors to be allocated more in under-reconstructed dynamic regions, achieving substantial improvement in reconstruction quality.
+Our anchor-based framework, coupled with the dynamic-aware anchor growing strategy, represents complex dynamic components in high-quality while addressing storage costs.
+We validate our approach through extensive experiments conducted on the N3DV (Li et al.
+2022) and Technicolor (Sabater et al.
+2017) datasets.
+Notably, our method outperforms all baselines by a large margin with practical storage overhead, as illustrated in Figure 1.
+Related Work
+Figure 2: Proposed 4D anchor-based framework.
+We begin with a sparse 4D grid of anchors, each defined by a unique 4D spatiotemporal position $\mathbf{p}$ and learnable offsets.
 Shared MLPs utilize these anchors to generate neural 4D Gaussians, capturing dynamic appearance changes with view direction $\mathbf{d}$.
-These 4D Gaussians are then projected to 3D Gaussians at specific timesteps for rendering, using a 3D Gaussian splatting pipeline to produce the final rendered image.
+These 4D Gaussians are then projected to 3D Gaussians at specific time $t_{r}$ for rendering, using a 3D Gaussian splatting pipeline to produce the final rendered image.
 We omit the $z$-axis for simplicity.
-2.1 Efficient 3D Gaussians
+Efficient 3D Gaussians.
 In recent years, 3D Gaussian Splatting (3DGS) has attracted significant attention for achieving real-time rendering by representing scenes with 3D Gaussian primitives and introducing a tile-based rasterizer.
 For photorealistic rendering quality, 3DGS requires a significant number of 3D Gaussians, which increases storage costs.
-To address this problem, [6, 7, 29] remove unnecessary Gaussians that do not harm rendering quality, [16, 19, 20] replace Gaussians or their parameters with efficient representations, and [19, 4, 33, 18] compress Gaussian parameters using existing compression techniques such as entropy coding and image compression.
-In this paper, we address eliminating excess Gaussians and improving efficiency in dynamic scenes with 4D Gaussians, which have been relatively unexplored compared to 3D Gaussians.
-2.2 Dynamic 3D Gaussians
-Two main approaches are proposed to extend 3DGS [8] into dynamic scene reconstruction.
-The first approach involves deforming 3D Gaussians along with temporal changes [34, 31, 2, 26, 15].
+To address this problem, previous works remove unnecessary Gaussians that do not harm rendering quality (Fan et al.
+2023; Girish, Gupta, and Shrivastava 2023), replace Gaussians or their parameters with efficient representations (Niedermayr, Stumpfegger, and Westermann 2024; Papantonakis et al.
+2024), and compress Gaussian parameters using existing compression techniques such as entropy coding and image compression (Niedermayr, Stumpfegger, and Westermann 2024; Chen et al.
+2025; Xie et al.
+2025).
+In this paper, we address compressing Gaussians in spatiotemporal space, which have been relatively unexplored compared to 3D Gaussians.
+Dynamic 3D Gaussians.
+Two main approaches are proposed to extend 3DGS (Kerbl et al.
+2023) into dynamic scene reconstruction.
+The first approach involves deforming 3D Gaussians along with temporal changes (Yang et al.
+2023; Wu et al.
+2023; Bae et al.
+2024; Shaw et al.
+2024; Jiawei et al.
+2024; Kwak et al.
+2025).
 These deformable 3D Gaussians offer the advantage of compact storage requirements but exhibit relatively slow rendering speeds and low visual quality.
-In contrast, the other approach directly employs 4D Gaussians in the spatio-temporal domain [35, 14].
-These approaches demonstrate superior visual quality and faster rendering speeds, but suffer from higher storage requirements.
-To address this, our method efficiently models temporal opacity and predicts Gaussian properties through neural features.
-2.3 Feature-based neural rendering
+In contrast, the other approach directly employs 4D Gaussians in the spatio-temporal domain.
+4DGS (Yang et al.
+2024) demonstrate superior visual quality and faster rendering speeds, but suffer from higher storage requirements.
+Although some works (Li et al.
+2023; Lee et al.
+2024a, c) improve storage efficiency using fewer Gaussians, they tend to focus on the holistic scene, thereby neglecting the quality of dynamic areas.
+To address this, we take an alternative approach: reducing storage while maintaining quality by preserving the Gaussian count.
+Feature-based neural rendering.
 Recently, a growing trend in scene reconstruction has been to integrate neural features as additional inputs to enhance model performance.
-For instance, there have been attempts to extract features from source views and utilize them for novel view synthesis, enabling few-view reconstruction [37, 3] or view interpolation through transformers [30, 22, 27].
-In 3DGS studies, [5] uses a hash grid instead of per-Gaussian SH coefficients, [14] renders features and decodes them into RGB colors via shallow MLP.
-[32] generates Gaussian attributes from a multi-level tri-plane.
-[16] predicts the attributes of local 3D Gaussians from anchor features.
-Our work focuses on efficient dynamic representations, which are relatively underexplored.
-We introduce a 4D anchor-based framework that includes dynamic neural velocity and temporal opacity derived from a generalized Gaussian distribution, considering real-world dynamics.
-3 Method
-We reconstruct dynamic 3D scenes with 4D Gaussians.
-Especially, our method builds 4D anchors which specify nearby 4D Gaussians (Sec.
-3.1).
-We encourage anchors to be closely placed at under-reconstructed dynamic regions (Sec.
-3.2).
-Furthermore, we design each 4D Gaussian to move along a line segment, and define the temporal opacity function which fits piecewise persistent period (Sec.
-3.3).
-3.1 Overview
+For instance, there have been attempts to extract features from source views and utilize them for novel view synthesis, enabling few-view reconstruction (Yu et al.
+2021; Chen et al.
+2021) or view interpolation through transformers (Wang et al.
+2021; Reizenstein et al.
+2021; T et al.
+2023).
+In 3DGS studies, Compact3DGS (Deng et al.
+2024) uses a hash grid instead of per-Gaussian SH coefficients and STG (Li et al.
+2023) renders features into RGB via shallow MLPs.
+Some works generate Gaussian attributes from a multi-level tri-plane (Wu and Tuytelaars 2024) and predict the attributes of local 3D Gaussians from anchor features (Lu et al.
+2024a).
+We introduce a 4D anchor-based framework that includes dynamic linear motion and temporal opacity derived from a generalized Gaussian distribution, considering real-world dynamics.
+Method
+We reconstruct dynamic 3D scenes through a sparse, grid-aligned 4D anchor grid.
+Each anchor holds a compressed feature vector that specifies nearby 4D Gaussians.
+The dynamic-aware anchor growing strategy encourages new anchors to be closely placed at under-reconstructed dynamic regions.
+Furthermore, we design each 4D Gaussian to move along a line segment, and define the temporal opacity function which fits a piecewise persistent period.
+Through these components, our framework employs a sufficient number of Gaussians to achieve high-quality reconstruction of dynamic regions, while the storage overhead is significantly reduced via compressed anchor features.
+4D Anchor-Based Framework
 The overview of our method is illustrated in Figure 2.
@@
 Corresponding 3D Gaussians are computed from these 4D Gaussians to render a frame at timestep $t$.
-We describe details of our method in the following paragraphs.
 Initializing 4D anchors.
-Similar with 3D Scaffold-GS [17], we utilize static point clouds to initialize 4D anchor points.
-We first obtain the static point cloud from multi-view frames at a certain timestep $t$ using Structure-from-Motion (SfM) [25].
-The 4D positions of anchors are initialized from a set of voxels $V\subset\mathbb{R}^{N\times 3}$ obtained from this point cloud:
-$p_{v}=(x_{v},y_{v},z_{v},t),\ \ \forall v\in V,$ (1)
+Similar to 3D Scaffold-GS (Lu et al.
+2024b), we utilize static point clouds to initialize 4D anchor points.
+We first obtain the static point cloud from multi-view frames at a certain timestep $t_{0}$ using Structure-from-Motion (SfM) (Schönberger and Frahm 2016).
+The 4D positions of anchors are initialized from a set of voxels $\mathbf{V}\in\mathbb{R}^{N\times 3}$ obtained from this point cloud:
+$\mathbf{p}=(x_{v},y_{v},z_{v},t_{0}),\ \ \forall v\in\mathbf{V},$ (1)
+$\mathbf{x}_{k}=\mathbf{p}+\Delta\mathbf{x}_{k},\ \ k\in\{1,2,...,K\},$ (2)
 where ($x_{v}$, $y_{v}$, $z_{v}$) is the center coordinates of the voxel $v$.
-Each anchor point also accompanies two learnable parameters: a feature vector $\mathbf{f}$, and 4D offsets $\mathbf{o}\in R^{K\times 3}$ for determining positions of $K$ neural 4D Gaussians.
-Beginning with 4D anchors initialized with the static point cloud, our method gradually increases anchors to dynamic regions.
-This is accomplished by our anchor growing strategy, which will be described in Sec.
-3.2.
+Each anchor point also accompanies two learnable parameters: a feature vector $\mathbf{f}$, and 4D offsets $\Delta\mathbf{x}_{k}\in\mathbb{R}^{4}$ for determining properties of $K$ neural 4D Gaussians.
+4D Gaussian position $\mathbf{x}$ is determined by anchor position and offset.
 Neural 4D Gaussians.
 We leverage shared MLPs and learnable parameters of the 4D anchors to generate $K$ neural 4D Gaussians from each anchor.
-Specifically, shared MLPs take the anchor feature $\mathbf{f}$ and yield properties of $K$ neural Gaussians, which include time-invariant opacity $\alpha\in\mathbb{R}$, quaternion $\mathbf{q}\in\mathbb{R}^{4}$ and scaling $\mathbf{s}\in\mathbb{R}^{3}$ for the covariance matrix, view-dependent color $\mathbf{c}\in\mathbb{R}^{3}$.
-Our shared MLPs also produce temporal scale $\sigma\in\mathbb{R}$ to compute our temporal opacity, and the neural velocity $\mathbf{v}\in\mathbb{R}^{3}$.
+Specifically, shared MLPs take the anchor feature $\mathbf{f}$ and yield properties of $K$ neural Gaussians, which include base opacity $\rho\in\mathbb{R}$, quaternion $\mathbf{q}\in\mathbb{R}^{4}$ and scaling $\mathbf{s}\in\mathbb{R}^{3}$ for the covariance matrix, view-dependent color $\mathbf{c}\in\mathbb{R}^{3}$.
+Our shared MLPs also produce a temporal scale $\sigma\in\mathbb{R}$ to compute our temporal opacity, and the neural velocity $\mathbf{u}\in\mathbb{R}^{3}$.
 Note that the color MLP additionally takes the view direction $\mathbf{d}\in\mathbb{R}^{3}$ as inputs for view-dependent modeling.
 Rendering neural Gaussians.
-To render the neural 4D Gaussians, we compute 3D Gaussian parameters at time $t$ from our time-varying properties.
-For the $k$-th Gaussian of the anchor $\mathbf{p}$, the opacity $\alpha_{k}(t)$ and the center $\mu_{k}(t)$ of the corresponding 3D Gaussian at time $t$ are derived as
-$\mu_{k}(t)=\mu_{k,1:3}+h(t,\mu_{k,4},\mathbf{v}),$ (2)
-$\alpha_{k}(t)=\alpha_{k}\cdot g(t,\mu_{k,4},\sigma_{k}),$ (3)
-where $\mu_{k}=\mathbf{p}+\mathbf{o}_{k}$ is the coordinates of the 4D Gaussian.
-$h(\cdot)$, $g(\cdot)$ model time-varying values of the positions and the opacity with the neural velocity and temporal opacity function, which will be further described in Sec.
-3.3.
-After deriving 3D Gaussians at time $t$, we utilize the existing efficient 3D Gaussian splatting pipeline [8] to render the frames.
-Same as 3D Scaffold-GS [17], Gaussians within the view frustum and having opacity higher than the threshold are passed to the renderer.
-Figure 3: Time-variant Gaussian position calculation.
-The 3D Gaussian position at rendering time is computed as the sum of the anchor position, the 4D Gaussian offset, and the temporal linear movement.
-The temporal linear movement is the product of the neural velocity and the time difference, allowing the Gaussian to traverse 3D space over time.
-3.2 Temporal coverage-aware anchor growing
+To render the neural 4D Gaussians, we compute 3D Gaussian parameters at time $t_{r}$ from our time-varying properties.
+For the $k$-th Gaussian of the anchor $\mathbf{p}$, the center $\mu_{k}\in\mathbb{R}^{3}$ and the opacity $\alpha_{k}\in\mathbb{R}$ of the corresponding 3D Gaussian at time $t_{r}$ are derived as
+$\mu_{k}=\mathbf{x}^{xyz}_{k}+h(t_{r},\mathbf{x}^{t}_{k},\mathbf{u}),$ (3)
+$\alpha_{k}=\rho_{k}\cdot g(t_{r},\mathbf{x}^{t}_{k},\sigma_{k}),$ (4)
+where $\mathbf{x}^{xyz}_{k}\in\mathbb{R}^{3}$ and $\mathbf{x}^{t}_{k}\in\mathbb{R}$ are the spatial and the temporal position of the 4D Gaussian.
+$h(\cdot)$, $g(\cdot)$ model time-varying values of the positions and the opacity with the neural velocity and temporal opacity function, which will be further described in the following section.
+After deriving 3D Gaussians at time $t_{r}$, we utilize the existing 3D Gaussian splatting pipeline (Kerbl et al.
+2023) to render the frames.
+Same as 3D Scaffold-GS (Lu et al.
+2024b), only Gaussians within the view frustum and having opacity higher than the threshold are passed to the renderer.
+Compact Parametrization of 4D Gaussians
+The properties of neural 4D Gaussians are predicted by a shallow shared MLP.
+Increasing the number of Gaussian properties complicates the optimization process.
+To this end, we propose a compact parametrization of 4D Gaussians that captures temporal changes as a set of linear segments.
+Our neural Gaussian design involves two key modelings.
+First, we represent time-varying spatial positions as linear motions.
+Second, we employ a generalized Gaussian function to model temporal opacity, which effectively captures sudden appearance changes with a single parameter.
+Linear motion.
+As described in Eq.
+(3), the time-varying 3D position of the 3D Gaussian $\mu_{k}\in\mathbb{R}^{3}$ is determined by the time-varying function $h(\cdot)$.
+We formulate the $h(\cdot)$ as a linear motion based on the neural velocity $\mathbf{u}$:
+$h(t,\mathbf{x}^{t}_{k},\mathbf{u})=(t-\mathbf{x}^{t}_{k})\mathbf{u}.$ (5)
+This formulation simplifies the temporal slicing of Gaussian positions proposed in 4DGS (Lee et al.
+2024a), using only 3 parameters per Gaussian.
+Despite its simplicity, a set of Gaussians with linear motion effectively captures the temporal dynamics of scene elements, particularly when combined with our modified temporal opacity.
+Figure 3: Illustration of the modified temporal opacity.
+To reconstruct the sudden and continuous appearance of an object, the univariate Gaussian function requires the sum of multiple Gaussians, while the generalized Gaussian achieves the same expressiveness with only a single Gaussian.
+Modified temporal opacity.
+Previous 4D Gaussian methods (Yang et al.
+2024; Li et al.
+2023) typically adopt the univariate Gaussian function to make Gaussians appear within a specific time range.
+However, this formulation lacks the capacity to model abrupt changes in real-world scene elements, limiting the expressiveness of individual Gaussians.
+We reformulate the temporal opacity function based on a generalized Gaussian distribution.
+This offers steeper derivatives at the beginning and the end, making it better suited to fit piecewise persistent periods.
+Our time-varying opacity $g(\cdot)$ described in Eq.
+(4) is formulated as
+$g(t,\mathbf{x}^{t}_{k},\sigma_{k})=exp({-(|t-\mathbf{x}^{t}_{k}|/\sigma_{k})^{\beta}}),$ (6)
+where $\beta$ is a tunable hyperparameter.
+In practice, we set the inverse sigma as a learnable parameter, as it leads to more stable training.
+We also model $\beta=2\beta^{\prime}$ and $\beta^{\prime}$ as a hyperparameter, which removes the $|\cdot|$ in the above equation.
+While the issue of temporal opacity is also explored in (Lee et al.
+2024a), we show that efficiently representing temporal coverage with fewer parameters is crucial for the performance of the anchor-based scheme.
+Dynamic-Aware Anchor Growing
 Growing new anchors to under-reconstructed dynamic regions is a critical factor for high-quality results.
 The direct application of previous anchor growing to the dynamic reconstruction fails to identify such regions, as it neglects the temporal coverage of the Gaussians.
-To this end, we propose a temporal coverage-aware anchor growing operation.
+The previous anchor growing strategy (Lu et al. 2024b) designed for static scenes gathers the gradients as a mean over $N$ iterations:
+$\nabla_{g}=\frac{\sum^{N}\|\nabla_{\text{2D}}\|}{N}.$ (7)
+With this strategy, dynamic regions appearing in a short period will have lower $\nabla_{g}$, as these regions will be penalized by the denominator $N$ regardless of their actual errors.
+To address this, we propose a dynamic-aware anchor growing operation.
 The core of our anchor growing is the computation of the accumulated gradients of each 4D Gaussian $\nabla_{g}$.
 We formulate $\nabla_{g}$ as a weighted sum of the gradient over $N$ iterations:
-$\nabla_{g}=\frac{\sum^{N}w(t,\sigma)\|\nabla_{\text{2D}}\|}{\sum^{N}w(t,\sigma)},$ (4)
-where $\nabla_{\text{2D}}$ is the 2D position gradient of the Gaussian. We define the weight term $w$ as a function of $\alpha(t)$ and $\sigma$, with $\gamma$ as a hyperparameter:
-$w(t,\sigma)=\alpha(t)(1/\sigma)^{\gamma}.$ (5)
-We collect the accumulated gradients of all Gaussians and then voxelize them. New anchors are placed at the centers of the voxels having gradients higher than a threshold.
-In contrast, the previous anchor growing strategy [17] designed for static scenes gathers the gradients as a mean over $N$ iterations:
-$\nabla_{g}=\frac{\sum^{N}\|\nabla_{\text{2D}}\|}{N}.$ (6)
-With this strategy, dynamic regions appearing in a short period will have lower $\nabla_{g}$, as these regions will be penalized by the denominator $N$ regardless of their actual errors.
-The first difference of our approach compared to the previous anchor growing strategy [17] is that the gradients are accumulated only when the Gaussian is activated.
-Our approach accurately gathers the gradients of 4D Gaussians placed at the regions where dynamic scene elements only appear in a short period.
-The second difference is that the temporal coverage $\sigma$ directly influences to the anchor growing operation.
-This makes regions having a short temporal coverage easier to grow.
-By modifying the accumulated gradients as a weighted sum, our approach effectively places new anchors to under-reconstructed dynamic regions.
-3.3 Modeling 4D Gaussians for compact anchors
-To represent dynamic scenes with fewer anchors, each neural Gaussian should be able to model the scene element in a sufficient time range.
-This necessitates neural 4D Gaussians to accurately model the real-world dynamics.
-We introduce two key formulations of 4D Gaussians that improve the capability of Gaussians of modeling real-world dynamics.
-We model local movements through the neural velocity, and we modify the temporal opacity based on a generalized Gaussian distribution.
-Neural velocity.
-In computer graphics, one common way to model the motion of an object is through a set of piecewise linear segments.
-Based on this concept, we propose the neural velocity to model the motions of scene elements with a set of 4D Gaussians having linear movements (see Figure 3).
-As described in Equation (2), the time-varying 3D position of the 3D Gaussian $\mu_{k}(t)\in\mathbb{R}^{3}$ is determined by the time-varying function $h(\cdot)$.
-We formulate the $h(\cdot)$ as the linear movement based on the neural velocity $\mathbf{v}$:
-$h(t,\mu_{k,4},\mathbf{v})=(t-\mu_{k,4})\mathbf{v}.$ (7)
-Note that our neural velocity simplifies the temporal slicing of 4D positions in 4DGS [10], without requiring excessive parameters for 4D covariances.
-Despite its simplicity, our neural velocity enables each 4D Gaussian to cover dynamic movements of the elements in a sufficient time range.
-Such capability reduces the number of neural 4D Gaussians and anchors as a result.
-Figure 4: Illustration of the modified temporal opacity.
-To reconstruct the sudden and continuous appearance of an object, the previous univariate Gaussian model requires the sum of multiple Gaussians, while the generalized Gaussian achieves the same expressiveness with only a single Gaussian.
-Modified temporal opacity.
-Previous 4D Gaussian methods [10, 14] adopt the univariate Gaussian function as temporal opacity to make Gaussians appear within a specific time range.
-However, this univariate Gaussian lacks the ability to model sudden changes in real-world scene elements, resulting in excessive Gaussians to model a single scene element.
-This issue is also discussed in the concurrent work [11], which proposes a simplified Gaussian mixture model as a solution.
-Therefore, we propose to modify the temporal opacity function to remove such redundant Gaussians (see Figure 4).
-We reformulate the temporal opacity function based on a generalized Gaussian distribution.
-This offers steeper derivatives at the beginning and the end, making it better suited to fit piecewise persistent periods.
-Our time-varying opacity $g(\cdot)$ in Equation (3) is formulated as
-$g(t,\mu_{k,4},\sigma_{k})=e^{-(|t-\mu_{k,4}|/\sigma_{k})^{\beta}}.$ (8)
-This formulation greatly reduces the number of 4D Gaussians and anchors to represent dynamic scenes while preserving the visual quality.
-In practice, we set the inverse sigma as a learnable parameter, as it leads to more stable training.
-We also model $\beta=2\beta^{\prime}$ and $\beta^{\prime}$ as a hyperparameter, which removes the ||$\cdot$|| ⋅ | in the above equation.
-3.4 Training and implementation details
-Table 1: Quantitative results on N3DV dataset.
-1 only includes the flame_salmon for storage and FPS [2].
-2 divides the scene into 50 frames and trains each using a separate model.
-3 uses COLMAP [25] for all 300 frames.
+$\nabla_{g}=\frac{\sum^{N}w(\alpha^{\prime},\sigma)\|\nabla_{\text{2D}}\|}{\sum^{N}w(\alpha^{\prime},\sigma)},$ (8)
+where $\nabla_{\text{2D}}$ is the 2D position gradient of the Gaussian, and $\alpha^{\prime}=g(t_{r},\mathbf{x}^{t},\sigma)$ is the time-variant component of the Gaussian opacity.
+We define the weight term $w$ as a function of $\alpha$ and $\sigma$, with $\gamma$ as a hyperparameter:
+$w(\alpha^{\prime},\sigma)=\alpha^{\prime}(1/\sigma)^{\gamma}.$ (9)
+We collect the accumulated gradients of all Gaussians and then voxelize them. New anchors are placed at the centers of the voxels having $\nabla_{\text{2D}}$ higher than each threshold.
+The dynamic-aware anchor growing strategy differs from the original anchor growing operation (Lu et al.
+2024a) in two key aspects.
+First, gradients are accumulated only when the Gaussian’s temporal opacity is greater than zero (i.e., when it is activated).
+It allows our method to accurately gather the gradients of 4D Gaussians placed at the regions where dynamic scene elements only appear in a short period.
+Second, the temporal coverage $\sigma$ directly influences to the anchor growing operation.
+This encourages regions with a short temporal coverage to receive stronger gradients, making it easier for anchors to grow in those areas.
+By accumulating gradients as a weighted sum, our method effectively places new anchors to under-reconstructed dynamic regions.
+Training and Implementation Details
+Dynamic region Full region Computational cost
 […]
-K-Planes [24] 1 31.22 0.947 0.090 309 MB 0.13
-MixVoxels-L [28] 1 30.81 0.933 0.095 512 MB 0.93
-HyperReel [1] 1 2 31.10 0.927 0.099 1362 MB 1.04
-4DGaussian [31] 30.71 0.935 0.056 59 MB 51.9
-E-D3DGS [2] 31.42 0.945 0.037 137 MB 19.9
-STG [14] 2 3 31.96 0.948 0.046 173 MB 125.9
-4DGS [35] 32.14 0.947 0.047 6194 MB 61.4
-Ours 32.03 0.947 0.041 134 MB 129.9
+4DGaussian (2023) 25.33 0.833 0.166 - 30.71 0.935 0.056 192 K 51.9 57 MB 50 m
+E-D3DGS (2024) 26.92 0.884 0.112 - 32.04 0.951 0.034 180 K 74.5 66 MB 1 h 52 m
+Grid4D (2024) 26.65 0.877 0.129 - 31.92 0.949 0.039 202 K 127.1 48 MB 1 h 20 m
+STG (2023) 25.84 0.860 0.127 44 K 31.24 0.941 0.051 434 K 125.9 63 MB 1 h 8 m
+4DGS (2024) 27.65 0.907 0.075 3306 K 32.14 0.947 0.047 3333 K 61.4 6194 MB 9 h 30 m
+4DGS (1GB) 26.70 0.877 0.123 557 K 31.59 0.943 0.052 581 K 152.7 1080 MB 3 h 37 m
+Ex4DGS (2024a) 26.33 0.874 0.121 52 K 32.01 0.947 0.048 268 K 91.9 115 MB 1 h 6 m
+Scaff-naive 24.79 0.811 0.199 206 K 31.21 0.942 0.053 732 K 159.7 83 MB 2 h 57 m
+Ours-light 27.50 0.902 0.076 181 K 31.54 0.944 0.045 314 K 148.2 90 MB 2 h 51 m
+Ours 28.86 0.927 0.054 533 K 32.03 0.947 0.041 775 K 129.9 149 MB 3 h 6 m
+Table 1: Quantitative results on the N3DV dataset. 4DGS refers to the result with 1GB storage, for fair comparisons.
+Figure 4: Qualitative comparisons on the N3DV dataset. Our method achieves high-fidelity in both static and dynamic regions.
 Training objective.
 Both learnable anchor parameters and MLP weights are jointly optimized through the rendering loss.
-We employ $\mathcal{L}_{1}$ with SSIM loss $\mathcal{L}_{\text{SSIM}}$ and volume regularization $\mathcal{L}_{\text{vol}}$ following the 3D Scaffold-GS [16].
+We employ $\mathcal{L}_{1}$ with SSIM loss $\mathcal{L}_{\text{SSIM}}$ and volume regularization $\mathcal{L}_{\text{vol}}$ following the 3D Scaffold-GS (Lu et al.
+2024a).
 The full training objective with weighting coefficients $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{vol}}=0.01$ is
@@
 Implementation details.
-Similar to 4DGS, our model is trained for 120K iterations with a single batch, which takes approximately 4 hours on a single NVIDIA A6000 GPU.
-We set $\beta=2$ and $\gamma=1$ for our modified temporal opacity model and temporal coverage-aware densification.
-For 4D voxel grid size, We use 0.001 as spatial grid size and use frame interval time as temporal grid size for all scenes.
-After training, we prune the unnecessary anchor points of which all Gaussians have opacity near zero.
-To improve rendering speed, we cache the time-invariant and view-invariant outputs of MLPs.
+Similar to 4DGS, our model is trained for 120K iterations with a single batch, which takes approximately 3 hours on a single NVIDIA A6000 GPU.
+We set $\beta=2$ and $\gamma=1$ for our modified temporal opacity model and dynamic-aware densification.
+For 4D voxel grid size, we use 0.001 as the spatial grid size and use the frame interval time as the temporal grid size for all scenes.
+We set $K=10$ following (Lu et al.
+2024a).
+After training, we prune the invalid anchor points of which all Gaussians have negative opacity.
+To improve inference speed, we cache the time-invariant and view-invariant outputs of MLPs.
 We provide more implementation details in the supplements.
-4 Experiments
+Dynamic region Full region Computational cost
+[…]
+4DGaussian (2023) 23.99 0.709 0.280 - 29.62 0.843 0.176 268 K 23.4 72 MB 31 m
+E-D3DGS (2024) 29.39 0.886 0.148 - 33.38 0.907 0.100 212 K 60.8 77 MB 2 h 34 m
+STG (2023) 28.35 0.869 0.174 71 K 33.33 0.912 0.096 132 K 149.0 30 MB 1 h 24 m
+4DGS (2024) 31.91 0.930 0.093 5597 K 33.30 0.910 0.095 5759 K 141.5 10699 MB 6 h 11 m
+4DGS (1GB) 29.18 0.890 0.144 538 K 28.72 0.856 0.166 591 K 156.6 1098 MB 3 h 36 m
+C3DGS (2024c) 27.21 0.843 0.209 87 K 32.52 0.901 0.110 122 K 178.7 18 MB 1 h 9 m
+Ex4DGS (2024a) 30.58 0.918 0.109 120 K 33.40 0.914 0.088 426 K 78.4 140 MB 1 h 53 m
+Scaff-naive 28.00 0.853 0.195 721 K 33.34 0.915 0.081 2056 K 127.0 129 MB 1 h 59 m
+Ours-light 30.95 0.916 0.117 186 K 33.97 0.923  0.074 480 K 131.4 108 MB 1 h 40 m
+Ours 31.94 0.932 0.093 1165 K 34.11 0.925 0.072 1272 K 110.8 278 MB 2 h 30 m
+Table 2: Quantitative results on Technicolor dataset. For fair comparison, 4DGS result of using 1GB storage is reported.
+Experiments
 In this section, we first evaluate our method through the comparisons with several state-of-the-art baselines for dynamic scene reconstruction.
@@
 Our code will be made publicly available.
-Datasets.   We evaluate our method on two representative real-world datasets: Neural 3D Video dataset (N3DV) [12] and Technicolor dataset [23].
+Datasets.  We evaluate our method on two representative real-world datasets, which are Neural 3D Video (N3DV) (Li et al. 2022) and Technicolor (Sabater et al. 2017).
   • Neural 3D Video dataset (N3DV) comprises 17 to 21 synchronized videos of six scenes, with each video containing 300 frames.
@@
 Baselines.
-We choose the following state-of-the-art competitors: 4DGaussian [31] and E-D3DGS [2] representing deformation-based methods, and 4DGS [35] and STG [14] representing 4D Gaussians.
-We reproduce them using their official code to report their performance.
-We also include representatives of NeRF-based methods: K-Planes [24], MixVoxels [28], and HyperReel [1].
-Results of these methods are brought from [2].
+We choose the following state-of-the-art competitors: 4DGaussian (Wu et al.
+2023), E-D3DGS (Bae et al.
+2024) and Grid4D (Jiawei et al.
+2024) representing deformation-based methods, and 4DGS (Yang et al.
+2024), STG (Li et al.
+2023), C3DGS (Lee et al.
+2024c) and Ex4DGS (Lee et al.
+2024a) representing 4D Gaussians.
+We reproduce them using the official code to report their performance.
+We also introduce Scaff-naive as an additional baseline.
+This is an anchor-based model without our neural Gaussian design and the anchor growing, which instead adopts linear motion and the temporal opacity modeling used in (Yang et al.
+2024).
+We present a variant of our model, Ours-light, which employs a larger voxel size to achieve a more reduced storage footprint.
+For fair comparisons, we report STG trained using 300 frames and initial SfM points derived from the first frames, same as other methods.
 Metrics.
-To assess the reconstruction quality, we measure peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and LPIPS [38] of the rendered images.
-We employ AlexNet [9] to compute LPIPS following the previous papers.
-To evaluate storage efficiency, we calculate the total size of output files, including MLP weights and parameters of Gaussians or anchors.
-We measure frames-per-second (FPS) by recording the wall-clock rendering time on the same machine, with a single NVIDIA A6000 GPU.
-Table 2: Quantitative results on Technicolor dataset. 1 refers to the values reported in [14]. 2 only includes the Painter for storage and FPS [2]. 3 uses COLMAP [25] for all 50 frames.
-[…]
-DyNeRF [12] 1 31.80 - 0.140 30 MB 0.02
-HyperReel [1] 2 32.32 0.899 0.118 289 MB 0.91
-4DGaussian [31] 29.62 0.844 0.176 51 MB 34.8
-E-D3DGS [2] 33.24 0.907 0.100 77 MB 36.5
-STG [14] 2 3 33.40 0.918 0.082 55 MB 112.7
-4DGS [35] 33.35 0.910 0.095 9106 MB 43.3
-Ours 33.97 0.923 0.074 98 MB 124.5
-Figure 5: Qualitative comparisons on the Neural 3D Video dataset.
-This figure provides qualitative comparisons on the Neural 3D Video dataset.
-White boxes indicate under-reconstructed areas.
-Our method achieves high fidelity in both dynamic and static regions.
-4.1 Reconstruction quality
+To assess the reconstruction quality, we measure peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and LPIPS (Zhang et al.
+2018) of the rendered images.
+To compute LPIPS, we follow (Bae et al.
+2024) and employ AlexNet (Krizhevsky, Sutskever, and Hinton 2012).
+To measure the visual quality in dynamic regions, we use a combined mask of Global-Median and Temporal-Difference (Li et al.
+2022), binarized with a threshold of 50.
+We assess storage efficiency by calculating the total size of the output files, including MLP weights, as well as Gaussian and anchor parameters.
+In addition to storage size, we also report the number of Gaussians in both dynamic and full regions.
+FPS and training time are measured on a single NVIDIA A6000 GPU, using the flame_salmon and Fabien scenes from N3DV and Technicolor, respectively.
+Experimental Results
 N3DV.
 Table 1 presents the quantitative results on the N3DV dataset.
-Our method outperforms all NeRF baselines in both visual quality and rendering speed.
-It also achieves competitive results in visual quality compared to state-of-the-art Gaussian Splatting-based methods, and offers the fastest rendering speed across all methods.
-Remarkably, our method significantly reduces storage costs by 97.8×97.8$\times$97.8 × compared to 4DGS [35], while achieving comparable or better performance in visual quality and rendering speed.
-It is worth noting that STG requires COLMAP for all frames, and train 6 models to obtain its results.
-We provide further comparisons with STG in the supplements.
-To further assess the performance, we deliver the qualitative comparisons with Gaussian Splatting baselines in Figure 5.
-Our method successfully reconstructs both static and dynamic regions.
-In contrast, other baselines often struggle to reconstruct dynamic regions (see the boxed regions).
-Deformation-based methods (4DGaussians, E-D3DGS) tend to produce blurry results, and the results of 4D Gaussian-based methods (4DGS, STG) often contain artifacts in dynamic regions.
-By effectively modeling dynamic regions with fewer anchors, our method delivers high-quality results in an efficient manner.
+Our model outperforms all baselines in terms of visual quality in dynamic regions, while also delivering competitive results in full regions.
+While our method employs more Gaussians to achieve high-quality results, it maintains efficient storage overhead and FPS, thanks to the anchor-based compression scheme.
+Our storage-efficient variant (Ours-light) achieves the second-best storage efficiency among 4D Gaussian methods, while also surpassing other efficient baselines in visual quality.
+Scaff-naive yields degraded results, indicating that the naïve extension of the scaffolding is insufficient to accurately reconstruct dynamic 3D scenes.
+4DGS often models static components through multiple dynamic Gaussians, resulting in an excessive number of Gaussians and storage costs.
+Although our method does not leverage additional techniques to model far-background areas (Li et al.
+2023; Yang et al.
+2024), it still exhibits plausible reconstruction quality in both static and full regions.
+The effectiveness of our method is also demonstrated in qualitative results.
+As shown in Figure 4, our model effectively represents dynamic regions with a sufficient number of Gaussians, thereby achieving high-quality dynamic reconstruction.
+In contrast, other efficient baselines struggle to capture complex scene dynamics (see the boxed regions).
 Please refer to the supplements for more visual comparisons, including uncropped results and videos.
-Figure 6: Qualitative comparisons on the Technicolor dataset.
-This figure provides qualitative comparisons on the Technicolor dataset.
-White boxes indicate under-reconstructed areas.
-Our method achieves high fidelity in both dynamic and static regions.
-Technicolor.
-We deliver the quantitative comparisons on the Technicolor dataset in Table 2.
-Our method surpasses all baselines in visual quality and rendering speed.
-We also achieve 85.1×85.1$\times$85.1 × storage reduction and $\times2$ improvement in FPS, compared to 4DGS.
-Although deformation-based methods offer compact storage usage, these methods require expensive computations for Gaussian deformations.
-This leads to the low FPS of these methods.
-Figure 6 presents the qualitative comparisons with Gaussian-based methods on the Technicolor dataset.
-Our method produces sharp and high-fidelity rendering results in both static and dynamic regions.
-On the other hand, 4DGaussians, E-D3DGS and STG show blurry results in dynamic regions, such as the boxed part in the Theater scene.
-E-D3DGS and STG often fail to reconstruct several details, e.g.
-the eye in the Theater scene.
-4.2 Analysis and ablation study
-To provide deeper insights of our main features, we conduct the ablations and analysis on the N3DV dataset. Please refer to the supplements for more detailed analysis and ablations on each component.
-Effects of each feature on quality and storage.
-We verify the effectiveness of our proposals through the ablations.
-We begin with the baseline model (Baseline) that does not employ none of our main components.
-Then we gradually add each component to the baseline and report its results.
-As shown in Table 3, our model with all components achieves the best results in visual quality with reasonable storage.
-Adding the temporal coverage-aware anchor growing operation improves visual quality but increases the anchors.
-The neural velocity reduces the number of anchors but brings degradation in visual quality.
-Finally, modifying the temporal opacity achieves the highest visual quality with fewer anchors.
-These results support that our anchor growing strategy successfully allocates new anchors to under-reconstructed dynamic regions, improving visual quality.
-Appropriate modeling of neural Gaussians reduces the number of anchors without sacrificing performance.
-Table 3: Ablation results on the main components.
-The abbreviations are as follows: “TA” refers to the temporal coverage-aware anchor growing, “NV” to the neural velocity and “MT” to the modified temporal opacity.
+Figure 5: Quality-storage trade-off.
+The results on the flame_steak scene of N3DV, along with the number of Gaussians, are reported.
+We increase the number of employed Gaussians of our model by adjusting the anchor grid size.
 […]
-31.20 0.053 207K 83 MB
-31.74 0.043  [^8] K 301 MB
-31.11 0.042 576K 195 MB
-31.82 0.043 793k 268 MB
-32.03 0.041 440K 149 MB
+Polynomial Ours 27.82 0.070 257 MB
+Linear 4DGS 27.93 0.065 195 MB
+Linear Ex4DGS 28.34 0.056 413 MB
+Linear Ours 25.77 0.153 180 MB
+Linear Ours 29.57 0.050 149 MB
+Table 3: Ablation results on the main components. “DA” refers to the dynamic-aware anchor growing, “Motion” and “Opacity” for each modeling of the neural Gaussian design.
 Figure 6: Anchor growing analysis.
@@
 Our anchor growing effectively accumulates gradients in under-reconstructed dynamic regions.
-Understanding anchor growing.
-To provide an in-depth understanding of the proposed anchor growing strategy, we present visual comparisons with previous anchor growing operations.
-We visualize the accumulated gradients $\nabla_{g}$ of both methods during training.
-We also compare resulting anchors and rendered images after training.
-Figure 7 presents visual comparisons with the previous anchor growing operation.
+Technicolor.
+We deliver the quantitative comparisons on the Technicolor dataset in Table 2.
+Our model achieves state-of-the-art performance across all visual quality evaluation metrics.
+This is further exemplified by our storage-efficient model, Ours-light.
+Despite its compact size of approximately 100 MB, Ours-light delivers competitive visual quality on par with 4DGS, highlighting the effectiveness of the proposed anchor-based scheme.
+Qualitative comparisons on the Technicolor dataset are provided in the supplements.
+Analysis and Ablation Study
+To provide deeper insights of our main features, we conduct the ablations and analysis in dynamic regions on N3DV.
+Quality-storage trade-off.
+We first analyze the quality-storage trade-off of our model by gradually increasing the number of employed Gaussians.
+This is controlled by adjusting the voxel size of the anchor grid.
+As presented in Figure 5, the performance of our model is improved with more Gaussians, confirming the necessity of using a sufficient quantity.
+Notably, our method outperforms other storage-efficient baselines when using similar or smaller storage costs.
+Thanks to the anchor-based scheme, our method employs $3.37\times$ or $7.0\times$ more Gaussians and achieves higher visual quality, yet still maintains a lower storage overhead.
+Understanding the anchor growing.
+We then ablate the proposed anchor growing strategy.
+As reported at the bottom of Table 3, the dynamic-aware anchor growing leads to noticeable improvement in visual quality.
+By properly allocating new anchors to under-reconstructed dynamic regions, our model accurately represents dynamic scenes with practical storage overhead.
+To further explore the effects of the anchor growing strategy, we provide visual comparisons in Figure 6.
 The previous method fails to accurately collect gradients in dynamic regions due to their short appearing periods.
-This leads to a lack of anchors in under-reconstructed regions, resulting in degraded visual quality.
+This leads to a lack of anchors in dynamic regions, resulting in degraded visual quality.
 In contrast, our method successfully accumulates higher gradients in these dynamic regions (see the red-boxed parts in Figure 6).
@@
 These results demonstrate that our anchor-growing strategy effectively captures under-reconstructed dynamic regions, playing a critical role in achieving high-quality reconstruction.
-Figure 8: Temporal opacity analysis.
-We visualize Gaussians that contributing to the rendering of a sample pixel in the sear_steak scene from the N3DV dataset.
-For clarity, we filter out Gaussians distant from the frying pan based on depth and image-plane probability.
-Our modified temporal opacity better aligns with real-world dynamics, allowing a single Gaussian to represent the element.
-Understanding temporal opacity.
+Effects of neural Gaussian design.
+We first evaluate our neural design by replacing each modeling with the formulation in previous works.
+Specifically, we replace the linear motion with the polynomial trajectory (Li et al.
+2023), and replace our modified temporal opacity with the one used in 4DGS (Yang et al.
+2024) and Ex4DGS (Lee et al.
+2024a).
+As shown in Table 3, our model with the proposed components achieves the best results.
+Our compact parametrization enhances the expressiveness of each Gaussian with minimal parameters, improving both visual quality and efficiency.
+Conclusion
+Our framework employs a sufficient number of Gaussians to capture complex dynamic regions, while addressing the storage overhead through the anchor-based compression.
+The dynamic-aware anchor growing and the neural Gaussian design lead to a substantial improvement.
+Experimental results support the validity of our method, achieving state-of-the-art visual quality and practical storage costs.
+Limitations and future work.
+Since our method is currently designed for multi-view video datasets, applying it to monocular videos can introduce additional challenges.
+Like other methods, our method still suffers from reconstructing elements that appear very shortly (1 or 2 frames).
+Resolving this challenge can be an interesting future work.
+Appendix
+Additional Comparison with STG
+Unlike our approach, which operates from sparse initial points, STG’s (Li et al.
+2023) original setting utilizes dense SfM point clouds from all timesteps and train the scene in 50-frame sequences.
+This original setting requires significant preprocessing for COLMAP (Schönberger and Frahm 2016) calculation and less densification process during training.
+Table S1 and Figure S1 show a qualitative and quantitative comparison between STG and our method, using the same number of iterations.
+While STG achieves higher scores when operating in its original dense setting, our method still demonstrates a superior score in visual quality.
+This indicates that our method successfully reconstructs fine-grained details that were omitted by the initial SfM points.
+Figure S1: Comparison between 50-frame STG and Ours.
+The single STG model trained on 300 frames of the N3DV dataset shows lower quality in the dynamic regions compared to Ours.
+The white boxes highlight under-reconstruction areas.
+[…]
+[…]
+[…]
+Ours 1 (300 frames) 120K 32.03 0.041
+Table S1: Comparison with STG on the N3DV dataset.
+Our model outperforms the average of six STGs trained on 50 frames and a single STG trained on 300 frames.
+$\ast$ Note that STG uses a batch size of 2 while our model uses a batch size of 1.
+Understanding Temporal Opacity
 We explore the effects of the temporal opacity by observing actual distributions.
@@
 We compare our temporal opacity with the previous one based on a univariate Gaussian function.
-As shown in Figure 8, the resulting Gaussian with our modified temporal opacity better fits the actual distributions of the object.
+As shown in Figure S2, the resulting Gaussian with our modified temporal opacity better fits the actual distributions of the object.
 This enables our method to represent a scene element with a single Gaussian.
@@
 As a result, our temporal opacity reduces the number of anchors while retaining visual quality, which leads to compact storage usage.
-5 Conclusion
-We proposed the 4D anchor-based method for dynamic scene reconstruction.
-Our anchor growing and neural 4D Gaussians modeling capture dynamic regions with an efficient number of anchors.
-Experimental results support the validity of our method, achieving compelling visual quality and FPS while significantly reducing storage consumption.
-Limitations and future work.
-Since our method is currently designed for multi-view video datasets, applying it on monocular videos can introduce additional challenges.
-Same as other methods, our method still suffer from reconstructing elements that appear very shortly (1 or 2 frames).
-Resolving this challenge can be an interesting future work.
-Supplementary Material
-Figure S1: Comparison between 300-frame STG and Ours.
-The single STG model trained on 300 frames of the N3DV dataset shows lower quality in the dynamic regions compared to Ours.
-The white boxes highlight under-reconstruction areas.
-Table S1: Comparison with STG on the N3DV dataset.
-Our model outperforms the average of six STGs trained on 50 frames and a single STG trained on 300 frames.
-$\ast$ Note that STG uses a batch size of 2 while our model uses a batch size of 1.
-[…]
-STG 6 (50 frames) 6$\times$30K 31.96 0.046
-1 (300 frames) 60K 31.25 0.053
-Ours 1 (300 frames) 120K 32.03 0.041
-Appendix A Additional comparison with STG
-As mentioned in the main paper, it is difficult to fairly compare with Spacetime Gaussians (STG) [14] due to the different experimental setups.
-Unlike our method, STG utilizes per-frame COLMAP to obtain the initial points and divide long video sequences into 50-frame chunks, training separate models for each chunk.
-This results in six separate models to report its results on N3DV.
-We additionally deliver the results of STG on N3DV, under the same experimental setup in Figure S1 and Table S1.
-While STG is effective on shorter video sequences, it often fails to accurately reconstruct dynamic regions on longer sequences.
-On the other hand, our method effectively represents dynamic regions on longer sequences.
-As a result, our method achieves better visual quality than STG under the same experimental setups.
-Appendix B Modified temporal opacity with 4DGS.
-We further validate the effectiveness of our modified temporal opacity by applying it to 4DGS [35].
-Consistent with the main paper, we visualize temporal opacity values at the same points within sear_steak scene of the N3DV dataset.
-$\beta=8$ is used for the generalized Gaussian equation.
-As shown in Figure S2, the previous univariate Gaussian function struggles to represent elements that suddenly appear and disappear.
-This limitation results in five Gaussians to model the same point.
-On the other hand, our modified temporal opacity effectively represents these dynamic changes, enabling the representation with only one Gaussian.
-Consequently, 4DGS with the modified temporal opacity achieves a storage reduction from 4.74 GB to 3.82 GB.
-Figure S2: Impact of modified temporal opacity with 4DGS.
-We apply our modified temporal opacity to 4DGS and analyze its behavior on the sear_steak scene from N3DV.
-The total temporal opacity distribution for the frying pan (blue point) is low at $t=20$ and $t=90$ while peaking at $t=60$, with rapid transitions in between.
-Each plot represents the temporal opacity of an individual Gaussian.
-Unlike conventional methods relying on Gaussian mixtures, our generalized Gaussian effectively captures this behavior using a single Gaussian component.
-Appendix C Additional ablation study
-Modeling 4D Gaussians without anchors.
-To analyze the effectiveness of our anchor-based reconstruction pipeline, we evaluate the performance of our model without the scaffold structure.
-Specifically, we replace all MLP outputs with explicit learnable Gaussian parameters.
-As shown in Table S2, removing the scaffold structure leads to a degradation in visual quality.
-The structured anchor representations offer more stable training and anchor growing, which significantly improves the performance of our method.
-Polynomial Gaussian trajectory.
-Our neural velocity represents the local dynamics of Gaussians as linear movements.
-Despite its simplicity, these linear movements sufficiently represent dynamics as 3D scene flows.
-We explore the effectiveness of the neural velocity by replacing it with a polynomial trajectory proposed in the previous method [14].
-We use $n_{p}=3,n_{q}=1$ as they use for their full model.
-The velocity MLP outputs coefficients of the polynomial equation instead of the neural 3D velocity.
-As shown in Table S2, this approach results in decreased performance and increased storage.
-Since the polynomial trajectory requires 12 parameters per Gaussian, a total of 120 parameters ($K=10$) as MLP outputs, it often complicates the training of our lightweight MLP and leads to degradation.
-Table S2: Additional model ablations. Quantitative results on different model designs on the N3DV dataset. Our full method utilizes scaffold structure and neural velocity.
-[…]
-velocity 30.11 0.060 393 MB
-polynomial 31.68 0.046 323 MB
-velocity 32.03 0.041 134 MB
-Appendix D Effects of hyperparameters
+Figure S2: Temporal opacity analysis.
+We visualize Gaussians contributing to the rendering of a sample pixel in the sear_steak scene from the N3DV dataset.
+For clarity, we filter out Gaussians distant from the frying pan based on depth and image-plane probability Our modified temporal opacity better aligns with real-world dynamics, enabling a single Gaussian to represent the element.
+Effects of Hyperparameters
 Figure S3: Effect of $\gamma$.
@@
 These new anchor points are primarily generated in under-reconstructed dynamic regions.
+$K=3$ $K=5$ $K=8$ $K=10$
+PSNR 27.40 28.63 29.09 29.20
+#Anchors 579K 585K 494K 476K
+Table S2: Results with various $K$.
 We investigate the impact of hyperparameters on the performance and their controllability. cook_spinach and flame_salmon scenes from N3DV are selected for analysis.
-Temporal coverage-aware anchor growing.
+Number of Gaussians per anchor.
+The parameter $K$ determines the number of Gaussians generated from an anchor feature.
+We observed changes in image quality and the number of anchors as $K$ varied.
+As $K$ increased, image quality improved while the number of anchors decreased.
+This is because a larger $K$ allows the same spatial region to be represented with higher density.
+Dynamic-aware anchor growing.
 The parameter $\gamma$ in Equation (9) controls the influence of temporal scale on densification.
@@
 We empirically observe that $\gamma=1$ provides a balanced reconstruction between static and dynamic regions.
-Modified temporal opacity.
+Temporal opacity.
 The parameter $\beta$ determines the steepness of our generalized Gaussian curve.
@@
 However, excessively high $\beta$ values (e.g., above 8) cause instability during training.
-Appendix E Visualization of anchor growing
+Visualization of Anchor Growing
 We visualize the anchor growing process in Figure S5.
@@
 As training progresses, anchors are created in the dynamic regions at each time step $t$ in under-reconstructed regions, improving image quality in these areas.
-Appendix F Additional method detail
+Additional Method Details
 Network architecture.
@@
 We employ a total of four MLPs, referred to as the Opacity MLP, Shape MLP, Color MLP, and Velocity MLP.
-The Opacity MLP outputs the time-invariant opacity $\alpha$ of the Gaussians.
+The Opacity MLP outputs the time-invariant opacity $\rho$ of the Gaussians.
 The Shape MLP produces the quaternion q and scaling s for covariance calculation, along with the temporal scale $\sigma$.
 The Color MLP takes the direction d concatenated with the anchor feature as input and outputs the view-dependent color c.
-The Velocity MLP computes the 3D velocity of the Gaussians.
-$\tanh$ for opacity $\alpha$, $\exp$ for scaling s and temporal scale $\sigma$, and sigmoid for color c are applied as activation function.
-Implmentation detail.
+The Velocity MLP computes the 3D velocity u of the Gaussians.
+$\tanh$ for opacity $\rho$, $\exp$ for scaling s and temporal scale $\sigma$, and sigmoid for color c are applied as activation function.
+Additional implementation detail.
 The temporal grid size is set to 0.0333 for the N3DV dataset and 0.02 for the Technicolor dataset.
 For initial points, we downsample it to fewer than 100,000 points.
-All other hyperparameters, including the learning rate and learning decay schedule, follow those of [16].
-Components such as the feature bank, LOD, and appearance features in [16] are excluded.
-Appendix G More results
-We report the quantitative results in Table S3-S4 and qualitative results in Figure S6-S7 of each scene of the N3DV and the Technicolor datasets, respectively.
+All other hyperparameters, including the learning rate and learning decay schedule, follow those of Scaffold-GS (Lu et al.
+2024a).
+Components such as the feature bank, level of detail (LOD), and appearance features are excluded.
+Results in Challenging Scenarios
+PSNR SSIM Train_time Storage
+4DGS 24.28 0.833 4 h 30 m 2359 MB
+Ours 28.82 0.913 2 h 25 m 114 MB
+Table S3: Results with limited views.
+Few shots.
+To evaluate the robustness of our method, we report the results in a 4 shot scenario, in which sparse 4 views are used for reconstruction.
+Results in a 4 shot scenario on the N3DV dataset are reported in Table S3.
+While our method was not specifically designed for few-shot scenarios, it exhibits a smaller performance drop compared to 4DGS.
+More Results
+We report the quantitative results in Table S4-S5 and qualitative results in Figure S7-S8 of each scene of the N3DV and the Technicolor datasets, respectively.
 We also provide video quality comparisons in HTML format.
 Please check the index.html file.
-Table S3: Per-scene quantitative results on the N3DV dataset.
-Model Metric coffee_martini cook_spinach cut_roasted_beef
+Figure S6: Qualitative comparisons on the Technicolor dataset.
+Dynamic region Full region
 […]
-4DGaussians 28.44 0.919 0.060 33.10 0.953 0.042 33.12 0.954 0.044
-E-D3DGS 29.10 0.931 0.042 32.96 0.956 0.034 33.57 0.958 0.034
-4DGS 28.63 0.918 0.071 33.54 0.956 0.039 34.18 0.959 0.038
-STG 28.26 0.915 0.074 33.07 0.958 0.038 33.38 0.959 0.039
-Ours 28.80 0.923 0.060 33.36 0.955 0.036 33.35 0.957 0.036
-Model Metric flame_salmon flame_steak sear_steak
+coffee_martini 4DGaussian 23.83 0.821 0.182 - 28.44 0.919 0.060 193K 58 MB
+E-D3DGS 26.05 0.901 0.096 - 29.10 0.931 0.042 263K 95 MB
+Grid4D 25.76 0.874 0.139 - 28.69 0.919 0.058 153K 36 MB
+STG 25.01 0.884 0.087 63K 27.45 0.912 0.070 481K 70 MB
+4DGS 26.90 0.920 0.065 4229K 28.63 0.918 0.071 4293K 7978 MB
+Ex4DGS 25.01 0.880 0.103 48K 28.79 0.918 0.067 340K 130 MB
+Ours 27.07 0.921 0.053 447K 28.80 0.923 0.060 729K 116 MB
+cook_spinach 4DGaussian 24.99 0.783 0.202 - 33.10 0.953 0.042 196K 58 MB
+E-D3DGS 26.08 0.824 0.140 - 32.96 0.956 0.034 140K 51 MB
+Grid4D 25.88 0.830 0.163 - 32.90 0.957 0.035 231K 55 MB
+STG 25.02 0.797 0.195 40K 32.25 0.948 0.049 430K 62 MB
+4DGS 27.32 0.879 0.088 3514K 33.54 0.956 0.039 3552K 6599 MB
+Ex4DGS 26.10 0.844 0.137 60K 33.22 0.955 0.041 247K 118 MB
+Ours 27.39 0.878 0.093 584K 33.36 0.955 0.036 782K 198 MB
+cut_roasted_beef 4DGaussian 27.19 0.834 0.183 - 33.12 0.954 0.044 184K 55 MB
+E-D3DGS 29.22 0.891 0.123 - 33.57 0.958 0.034 143K 52 MB
+Grid4D 28.74 0.886 0.133 - 33.61 0.958 0.034 229K 54 MB
+STG 27.87 0.864 0.150 38K 32.73 0.951 0.048 383K 56 MB
+4DGS 31.21 0.936 0.060 3010K 34.18 0.959 0.038 3047K 5661 MB
+Ex4DGS 29.56 0.904 0.099 64K 33.72 0.957 0.039 249K 123 MB
+Ours 31.80 0.953 0.043 745K 33.35 0.957 0.036 916K 170 MB
+flame_salmon 4DGaussian 21.98 0.832 0.163 - 28.80 0.926 0.060 197K 58 MB
+E-D3DGS 24.29 0.890 0.104 - 29.61 0.936 0.038 264K 96 MB
+Grid4D 24.52 0.888 0.113 - 29.86 0.930 0.051 158K 37 MB
+STG 22.09 0.848 0.120 62K 28.27 0.916 0.065 562K 81 MB
+4DGS 24.43 0.900 0.078 4774K 29.25 0.929 0.057 4782K 8886 MB
+Ex4DGS 22.75 0.860 0.128 49K 28.78 0.924 0.064 330K 128 MB
+Ours 24.82 0.911 0.059 473K 29.20 0.928 0.054 896K 162 MB
+flame_steak 4DGaussian 25.43 0.855 0.134 - 33.55 0.961 0.032 186K 56 MB
+E-D3DGS 26.54 0.891 0.101 - 33.57 0.964 0.028 135K 50 MB
+Grid4D 25.77 0.880 0.117 - 32.98 0.963 0.029 214K 50 MB
+STG 25.95 0.868 0.122 31K 33.35 0.958 0.037 373K 54 MB
+4DGS 26.74 0.893 0.081 2439K 33.90 0.961 0.037 2451K 4553 MB
+Ex4DGS 27.03 0.905 0.077 47K 33.90 0.962 0.032 226K 101 MB
+Ours 27.84 0.922 0.049 500K 33.43 0.962 0.030 734K 142 MB
+sear_steak 4DGaussian 28.59 0.876 0.127 - 34.02 0.963 0.031 198K 59 MB
+E-D3DGS 29.36 0.906 0.107 - 33.45 0.963 0.030 135K 49 MB
+Grid4D 29.27 0.905 0.109 - 33.49 0.965 0.028 228K 54 MB
+STG 29.09 0.901 0.087 32K 33.40 0.960 0.034 371K 54 MB
+4DGS 29.31 0.913 0.082 1866K 33.37 0.960 0.040 1880K 3492 MB
+Ex4DGS 30.52 0.927 0.063 41K 33.68 0.960 0.034 221K 94 MB
+Ours 31.44 0.945 0.045 448K 33.52 0.960 0.030 595K 109 MB
+Table S4: Quantitative results on the N3DV dataset.
+Dynamic region Full region
 […]
-4DGaussians 28.80 0.926 0.060 33.55 0.961 0.032 34.02 0.963 0.031
-E-D3DGS 29.61 0.936 0.038 33.57 0.964 0.028 33.45 0.963 0.030
-4DGS 29.25 0.929 0.057 33.90 0.961 0.037 33.37 0.960 0.040
-STG 29.44 0.925 0.067 33.82 0.965 0.031 33.77 0.966 0.031
-Ours 29.20 0.928 0.054 33.73 0.960 0.030 33.73 0.960 0.031
-Table S4: Per-scene quantitative results on the Technicolor dataset.
-Model Metric Birthday Fabien Painter
-[…]
-4DGaussians 28.03 0.862 0.156 33.36 0.865 0.185 34.52 0.899 0.138
-E-D3DGS 33.34 0.951 0.037 34.45 0.875 0.147 36.63 0.923 0.097
-4DGS 31.81 0.937 0.040 35.02 0.884 0.144 35.71 0.923 0.104
-STG 31.88 0.947 0.028 35.79 0.900 0.115 35.53 0.925 0.095
-Ours 32.86 0.952 0.024 35.32 0.889 0.140 37.35 0.941 0.069
-Model Metric Theater Train
-[…]
-4DGaussians 28.67 0.835 0.195 23.54 0.756 0.206
-E-D3DGS 30.64 0.864 0.147 31.84 0.922 0.074
-4DGS 31.94 0.875 0.143 32.25 0.932 0.046
-STG 31.25 0.880 0.133 32.58 0.938 0.039
-Ours 31.65 0.891 0.106 32.69 0.941 0.032
-Figure S6: Qualitative comparisons on the N3DV dataset.
-Figure S7: Qualitative comparisons on the Technicolor dataset.
+Birthday 4DGaussian 25.14 0.838 0.150 - 28.03 0.862 0.156 297K 79 MB
+E-D3DGS 29.10 0.948 0.041 - 33.34 0.951 0.037 231K 84 MB
+STG 28.98 0.950 0.037 123K 32.20 0.947 0.029 302K 44 MB
+C3DGS 28.03 0.933 0.054 126K 31.37 0.938 0.039 177K 26 MB
+4DGS 29.27 0.953 0.032 7233K 31.72 0.936 0.042 7366K 13684 MB
+Ex4DGS 29.07 0.949 0.033 169K 32.25 0.942 0.031 460K 162 MB
+Ours 29.51 0.956 0.026 395K 32.86 0.952 0.024 623K 183 MB
+Fabien 4DGaussian 28.54 0.807 0.283 - 33.36 0.865 0.185 224K 61 MB
+E-D3DGS 30.23 0.859 0.214 - 34.45 0.875 0.147 173K 63 MB
+STG 29.90 0.848 0.240 26K 34.69 0.876 0.166 72K 10 MB
+C3DGS 29.54 0.836 0.260 30K 34.36 0.872 0.174 44K 6 MB
+4DGS 31.44 0.892 0.170 2304K 35.02 0.884 0.144 2327K 4322 MB
+Ex4DGS 31.33 0.890 0.155 116K 34.98 0.889 0.131 202K 83 MB
+Ours 31.44 0.896 0.151 380K 35.15 0.895 0.124 398K 136 MB
+Painter 4DGaussian 27.31 0.819 0.187 - 34.52 0.899 0.138 184K 51 MB
+E-D3DGS 31.60 0.916 0.124 - 36.63 0.923 0.097 206K 75 MB
+STG 32.10 0.925 0.120 61K 36.66 0.923 0.097 140K 20 MB
+C3DGS 30.39 0.903 0.145 61K 35.55 0.913 0.113 75K 11 MB
+4DGS 33.51 0.946 0.097 4945K 35.71 0.923 0.104 5015K 9316 MB
+Ex4DGS 33.33 0.945 0.096 151K 36.62 0.932 0.091 255K 106 MB
+Ours 35.19 0.960 0.075 1729K 37.05 0.939 0.073 1834K 465 MB
+Train 4DGaussian 17.11 0.470 0.398 - 23.54 0.756 0.206 332K 88 MB
+E-D3DGS 28.48 0.914 0.068 - 31.84 0.922 0.074 216K 78 MB
+STG 26.56 0.896 0.110 101K 32.20 0.935 0.044 314K 46 MB
+C3DGS 25.23 0.855 0.174 138K 30.79 0.910 0.071 193K 28 MB
+4DGS 33.09 0.956 0.034 6468K 32.15 0.930 0.047 6980K 12967 MB
+Ex4DGS 29.97 0.943 0.058 48K 31.40 0.927 0.055 816K 216 MB
+Ours 32.42 0.964 0.026 2339K 32.86 0.946 0.032 2433K 275 MB
+Theater 4DGaussian 21.86 0.612 0.381 - 28.67 0.835 0.195 302K 80 MB
+E-D3DGS 27.56 0.794 0.291 - 30.64 0.864 0.147 237K 86 MB
+STG 24.19 0.727 0.360 45K 30.92 0.878 0.143 188K 27 MB
+C3DGS 22.85 0.690 0.411 81K 30.54 0.872 0.155 123K 18 MB
+4DGS 32.25 0.904 0.132 7036K 31.90 0.877 0.136 7110K 13208 MB
+Ex4DGS 29.21 0.862 0.204 119K 31.74 0.879 0.131 399K 132 MB
+Ours 31.12 0.885 0.188 980K 32.65 0.893 0.106 1073K 332 MB
+Table S5: Quantitative results on the Technicolor dataset.
+Figure S7: Qualitative comparisons on the N3DV dataset with full region.
+Figure S8: Qualitative comparisons on the Technicolor dataset with full region.